### v2

# Visual Decoding and Reconstruction via EEG Embeddings with Guided Diffusion

Abstract
How to decode human vision through neural signals has attracted a long-standing interest in neuroscience and machine learning. Modern contrastive learning and generative models improved the performance of fMRI-based visual decoding and reconstruction. However, the high cost and low temporal resolution of fMRI limit their applications in brain-computer interfaces (BCIs), prompting a high need for EEG-based visual reconstruction. In this study, we present an EEG-based visual reconstruction framework. It consists of a plug-and-play EEG encoder called the Adaptive Thinking Mapper (ATM), which is aligned with image embeddings, and a two-stage EEG guidance image generator that first transforms EEG features into image priors and then reconstructs the visual stimuli with a pre-trained image generator. Our approach allows EEG embeddings to achieve superior performance in image classification and retrieval tasks. Our two-stage image generation strategy vividly reconstructs images seen by humans. Furthermore, we analyzed the impact of signals from different time windows and brain regions on decoding and reconstruction. The versatility of our framework is demonstrated in the magnetoencephalogram (MEG) data modality. We report that EEG-based visual decoding achieves SOTA performance, highlighting the portability, low cost, and high temporal resolution of EEG, enabling a wide range of BCI applications. Our code is available at https://github.com/dongyangli-del/EEG_Image_decode.
Keywords:
EEG Visual reconstruction Neural alignment MEG
[…]
Figure 1: EEG/MEG-based visual retrieval, classification, and reconstruction tasks. (a) Overview of three visual decoding tasks using EEG/MEG data under natural image stimuli. (b) Comparisons of nine encoders on the THINGS-EEG dataset, including within-subject and cross-subject performance. (c) Comparisons on the THINGS-MEG dataset, similar to (b). Our ATM-S encoder achieves the highest performance compared to other competing encoders in EEG/MEG-based visual decoding tasks.
1 Introduction
A key technical challenge in brain-computer interfaces (BCIs) is to decode/reconstruct the visual world seen by humans through non-invasive brain recordings, such as functional magnetic resonance imaging (fMRI), magnetoencephalography (MEG) or electroencephalography (EEG). These highly dynamic brain activities reflect human perception of the visual world, which is influenced by properties of the external visual stimulus, our internal states, emotions and even personal experiences. Thus, visual decoding and reconstruction based on neural signals can uncover how the human brain processes and interprets natural visual stimuli, as well as promote non-invasive BCI applications.
Figure 2: EEG/MEG-based visual decoding and generation framework. The EEG encoder is designed as a flexible replacement component. After aligning with image features, the EEG features are used for zero-shot retrieval and classification tasks, and the reconstructed images are obtained through a two-stage generator.
Contrastive learning and generative models have greatly advanced fMRI-based visual decoding in both decoding tasks (e.g., image classification and retrieval) and generative tasks (e.g., image reconstruction). By combining pre-trained visual models, existing fMRI decoding models can learn highly-refined feature embeddings in limited data [28, 40]. Using these embedded fMRI features, generative models such as diffusion models can reconstruct the image one is seeing [40, 12]. However, despite many advances in fMRI-based visual decoding, fMRI equipment is unportable, expensive, and difficult to operate, largely limiting its application in BCIs. Alternatively, EEG is portable, cheap, and universal, facilitating a wide range of BCI applications. EEG has higher temporal resolution and can effectively capture rapid changes in brain activity when processing complex, dynamic visual stimuli.
However, EEG has long been considered incomparable to fMRI in natural image decoding/reconstruction tasks, as EEG suffers from low signal-to-noise ratio, low spatial resolution, and large inter-subject variability. Recent advances in multimodal alignment have made MEG/EEG visual decoding possible, although the performance is still inferior to fMRI [8, 46, 16]. Yohann Benchetrit et al. used the CLIP model to extract the latent representation of the image and trained the MEG encoder to align it with the image representation extracted by CLIP. It achieved excellent retrieval and reconstruction performance on MEG and fMRI datasets, demonstrating the potential for real-time visual decoding and reconstruction using EEG/MEG signals. Recently, Song et al. [46] used an EEG encoder based on ShallowNet [37] and performed representation alignment through contrastive learning, achieving excellent decoding performance on the THING-EEG dataset [14]. These two studies provide preliminary evidence of the potential of EEG/MEG-based visual decoding. However, there is a significant gap in their performance compared to the fMRI-level performance. This gap is largely caused by weaknesses in its EEG encoder, which prevents EEG embeddings from effectively aligning with image embeddings.
To fill in the gap, we develop an EEG/MEG-based visual decoding framework, including a novel EEG encoder, called Adaptive Thinking Mapper (ATM), and a two-stage image generation strategy. We survey existing EEG encoder modules, such as EEGNetV4  [25], ShallowNet [37], and Conformer  [44], and demonstrate through extensive comparative experiments and ablation studies that our method achieves state-of-the-art (SOTA) performance on image classification, retrieval, and generation tasks. Our work has three main contributions:
  • We employ the attention module and a spatiotemporal convolution module in EEG encoder, which can be used in a plug-and-play manner, to extract representations in real-time for both EEG data and MEG data.
  • Our brain encoder is trained with self-supervised contrastive learning framework, achieving SOTA performance on EEG and MEG datasets from natural image decoding tasks.
  • We present a two-stage image generation strategy to facilitate the adjustment of the distribution from the original EEG data to the corresponding images. It employs a prior diffusion model that is conditioned on the EEG embeddings to generate the image embeddings, followed an enhanced Stable Diffusion module to generate images with EEG priors.
4
Figure 3: The structure of EEG encoder ATM. According to the position of EEG channels and the spatiotemporal property of EEG signals, we design the position encoding and temporospatial encoding in ATM.
2 Method
To learn high-quality latent representations of EEG data, it is crucial to consider the spatial position of EEG channels and the spatiotemporal properties of EEG signals. Let $T$ represent the length of the time window of the data, $C$ the number of EEG channels, and $N$ the total number of data samples. Our objective is to derive EEG embeddings $Z_{E}=f(E)\in\mathbb{R}^{N\times F}$ from the brain activity data $E\in\mathbb{R}^{N\times C\times T}$, where $f$ is the EEG encoder and $F$ is the feature dimension of the embeddings. Concurrently, we use the CLIP model to extract image embeddings $Z_{I}\in\mathbb{R}^{N\times F}$ from images $I$. Our goal is to effectively align the EEG representation with the image representation, as illustrated in Fig. 2. In the training phase, the EEG encoder is trained with EEG and image pairs using a contrastive learning framework. In the inference phase, the EEG embeddings from the trained plug-and-play EEG encoder can be used for a variety of tasks, including EEG-based image classification, retrieval, and EEG-guided image generation.
2.1 ATM for EEG embedding
We develop an EEG encoder, called Adaptive Thinking Mapper (ATM), for aligning the original EEG signals to its feature representation space (Fig. 3). ATM is based on the Transformer Encoder and spatiotemporal convolution architecture. Specifically, we utilize a one-dimensional linear layer to project the input one-dimensional EEG data to the embedding dimension required by the Transformer model. These inputs are processed through a self-attention module to integrate the embeddings of input data and positional encoding. We first divide the input into fixed-size patches, with each image patch containing a fixed number of data points. Then, each patch is mapped to a high-dimensional embedding space through a linear layer. Subsequently, through the temporal aggregation process, we obtain the processed embedding representations. Notably, ATM addresses the inadequacies of other modules in modeling capabilities on the temporal scale through the Input Layer. For it has been reported that the spatiotemporal convolution module with a large convolution kernel is an effective way to represent EEG data with a small number of parameters [46], we use a similar spatiotemporal convolution (STConv)-based feature extraction module. The difference is our STConv Module is plug-and-play and can be flexibly replaced with different types of spatiotemporal convolution components as needed to adapt to various EEG/MEG datasets. Finally, MLP projection layer consists of $M$ simple residual components and fully connected layers, with LayerNorm applied in the output to ensure the stability of training.
2.2 Image Embedding
Many previous studies have explored various training strategies to train deep neural networks for image embedding, such as VGG-19 and ResNet trained with supervised learning, CLIP trained with contrastive learning, and VAEs with self-supervised learning [49, 4, 46, 16]. They have reported that CLIP models pre-trained using the Vision Transformer (ViT) architecture perform better in a range of downstream tasks, including image decoding and reconstruction, compared to models trained using supervised learning methods (such as VGG, ResNet) and self-supervised VAE frameworks. Thus, in this study, we use CLIP for image embedding and align the CLIP’s output image representation, denoted as $Z_{I}\in\mathbb{R}^{N\times 1024}$, with the EEG embeddings. Before formal training, all images undergo standard preprocessing  [34].
2.3 EEG guidance image generation
In this study, we present a two-stage pipeline for generating images that serve as visual stimuli for EEG recordings, as shown in the bottom right of Fig. 2. In the left of Fig. 3 we have obtained the EEG embeddings $z_{E}$ for each image by the EEG encoder ATM. Now our goal is to use these EEG embeddings to generate the corresponding images. The joint distribution of images, EEG embeddings, and image embeddings can be expressed as $p(I,z_{E},z_{I})=p(z_{I}|z_{E})p(I|z_{I})$, corresponding to the prior diffusion and CLIP-guided generation, respectively. In Stage I, we first focus on the prior diffusion stage. Inspired by DALL-E 2 [35] and Mind’s Eyes [40], we train a diffusion model conditioned on the EEG embeddings $\hat{Z}_{E}$ to learn the distribution of CLIP embeddings $p(z_{I}|z_{E})$. In this stage, we construct a lightweight U-Net: $\epsilon_{\text{prior}}(z^{t}_{I},t,z_{E})$, where $z^{t}_{I}$ represents the noisy CLIP embedding at diffusion time step $t$. We train the prior diffusion model using EEG and CLIP embeddings. Through this diffusion model, we can generate corresponding CLIP embeddings $z_{I}$ from EEG embeddings as a prior for stage II. In Stage II, we employ the pre-trained SDXL [33] and IP-Adapter [59] models to model the generator $p(I|z_{I})$, thereby sampling image $I$ according to $z_{I}$. This stage mainly focuses on converting CLIP embeddings into corresponding images. Further details are provided in Appendix C.
2.4 Loss Function
Following the methodology outlined by Benchetrit et al. [4], we adopt a dual approach to loss functions, serving distinct objectives. For the classification and retrieval tasks, we only utilize the CLIP loss, which is inspired by the contrastive learning approach described in Radford et al. [34]. This loss function aids in aligning the EEG data $E$ with corresponding image data $I$, thereby facilitating the identification of EEG-image pairs and maximizing the EEG classification boundaries. For the generation tasks, besides the CLIP loss, we add a Mean Squared Error (MSE) loss to facilitate learning in regression scenarios. Thus the overall loss function for our model is a combination of these two distinct loss types, expressed as:
$Loss=\lambda\cdot L_{CLIP}+(1-\lambda)\cdot L_{MSE}$
Here, $\lambda$ is a hyperparameter that balances the contribution of each loss type.
Figure 4: EEG-based image classification. (a) The paradigm of EEG-based image classification. (b) Average accuracy across different methods in the subjects.
Figure 5: EEG-based image retrieval. (a) The paradigm of EEG-based image retrieval. (b) Average accuracy across different methods in the subjects. (c) Images showing the top-5 accuracy in EEG-image retrieval tasks. See Appendix E for additional images results. (d) Representation similarity analysis. Clustering of image features and EEG features using a k-means algorithm with k=5 is performed, followed by the calculation of similarity between clusters. See Appendix D for additional Representational similarity results.
3 Experiments
3.1 Training and computational considerations
We conducted our experiments on the THINGS-EEG dataset’s training set [14, 16]. To verify the versatility of ATM for embedding electrophysiological data, we tested it on MEG data modality using the THINGS-MEG dataset [18]. We used the Adam optimizer [24] to train the across-subject model on a set of approximately 496,200 samples, and the within-subject model on a set of about 66,160 samples, with a learning rate of $3\times 10^{-4}$ and batch sizes of 16 and 1024. Our initial temperature parameter was set to 0.07. We tested on the zero-shot test dataset at the end of each training epoch during the training process. For fairness, all models’ hyperparameters were kept consistent. In our study, we compared the performance of different encoders on the within-subject test set and cross-subject (leave-one-subject-out) test set (see Appendix F).
3.2 EEG Decoding performance
The plug-and-play ATM can obtain the EEG embedding for the classification task using a simple cosine similarity measurement. We output the category with the highest cosine similarity (Fig. 4a). More details of the EEG-based image classification are in Appendix B. Fig. 4b presents the accuracy of EEG classification for each subject, and Fig. 4c shows our method outperforms others.
Here, we test the effectiveness of ATM-extracted EEG embeddings in the image retrieval task. Fig. 5a shows the image retrieval process. We calculate the cosine similarity between the extracted EEG embeddings and the CLIP embeddings of the image dataset (with 200 images), and output the image with the highest similarity as the retrieved image. Fig. 5b shows the Top-5 retrieved images corresponding to the real visual stimuli seen by subjects. To be a fair comparison, here we presented the images that the previous method failed to retrieve in Top-1 for testing [46]. Compared with the previous model, the Top-1 accuracy of our model is significantly improved, and the Top-5 images all maintain a high degree of similarity with the original images. Fig. 5c shows our method outperforms others.
Figure 6: EEG guidance image generation. (a) The paradigm of EEG guidance image generation. (b) The similarity between random visual objects and the EEG embeddings, and the similarity between generated visual objects and the target EEG embeddings. (c) Comparison between the original image and the image generated using the corresponding EEG data. (see Appendix B & C for details). (d) The similarity between visual objects and target EEG embeddings as the guidance scale changes, and the diversity of visual objects as the guidance scale changes. See Appendix E for additional results.
Ablation study on ATM
We systematically deconstructed and analyzed each layer of our EEG encoder module. We conducted an ablation study for each component in ATM (i.e., the MLP projector, the adaptive convolution module, the subject-wise linear, and the spatial attention block). Notably, ATM is a flexible, plug-and-play module, which can be easily replaced with any form of convolutional component or even omitted. Appendix B.3 showed that the adaptive convolutional components significantly enhanced encoder performance, while the subject-wise linear layer reduced the model’s training consumption and substantially improved the existing components’ capabilities. Surprisingly, our spatial attention module based on sine-cosine positional encoding improved encoder performance only when used alone, but the benefits were not as significant when combined with the subject-wise linear layer.
3.3 Image Generation performance
Using the image generation task, we verify the effectiveness of ATM’s EEG embedding and our two-stage generation model. Fig. 6a shows the process of generating images under the guidance of EEG embedding and evaluating the quality of the generated images. For one test image, we put EEG embedding of the subject viewing the image into the two-stage generator to guide the generation of the image. To evaluate the generation performance, we conducted an image retrieval task. Specifically, we extract the CLIP embedding of the generated images and calculate the similarity between the CLIP embeddings of all images in the image dataset to retrieve the generated image.
Figure 7: Examples of EEG-guided visual reconstruction. From top to bottom, we exhibit the best, median, and worst 12 generated images, respectively. We show the images subjects seen and the generated images by our two-stage image generator. See Appendix E for additional results.
Fig. 6b shows the generation effect of the sample image. Fig. 6c shows the retrieval accuracy using generated images. The generated images have high semantic similarity with the seen images and have good diversity in low-level visual features, which can be manipulated by the guidance scale hyperparameter (Fig. 6d). Surprisingly, the retrieval accuracy using generated images is even higher than that using EEG embeddings, implying the benefit of our two-stage visual reconstruction for image retrieval (See Appendix C for Figure A2). We also report the decoding and reconstruction performance for EEG, MEG, and fMRI across various metrics in the Appendix D.
We visualize the best, medium and worst generated images in Fig. 7. We randomly selected the EEG data of a subject viewing 100 images, and extracted EEG embeddings to guide image generation. By calculating the cosine similarity of the CLIP embedding between the generated image and the original image, we found 12 images each with the best, medium and worst generation effects. It can be seen that in the best group, the generated image is not only highly consistent with the semantics of the original image, but also well retains the low-level visual features. in the medium group, the generated image maintains the semantic features of the original image, and the low-level visual features are well preserved. Visual features were altered. in the worst group, both semantic features and low-level visual features were altered.
3.4 Temporal analysis
To investigate the effects of EEG time window on visual decoding accuracy, we calculated the average top-1 classification accuracy for two different time windows: [0, $t$], including the entire period from the onset of visual stimuli to time point $t$, and [$t$-100, $t$], only including the data 100ms before time point $t$. We compared the accuracy with a randomly selected baseline (0.5% chance level) to test non-random predictive performance (Fig. 8). Our results show that within 500ms after visual stimulus presentation, the EEG signal decoding accuracy reaches an upper limit of about 30%, after which the accuracy no longer improves (Fig. 8a). The MEG decoding shows a similar profile as the time window expands (Fig. 8b). We exhibit the generated images under different EEG time windows, [0, $t$] in Fig. 8c. The similarity between the generated images and the original images is low when the time window is less than 150ms, then this similarity gradually increase as the time window expands. After 500 milliseconds, EEG-guided image generation can reliably reveal the semantics of the images seen, such as aircraft carriers, modems, jelly beans, et al. Interestingly, we find differences in the optimal reconstruction time windows for different categories of images, for example, jelly beans (200ms) are faster than aircraft carrier (500ms), implying that the human brain may process different visual objects at different speeds. This finding highlights the advantage of EEG’s high temporal resolution in studying fast visual processing compared with the lower temporal resolution of fMRI.
Figure 8: Effects of different EEG/MEG time windows on EEG-guided visual retrieval and reconstruction. (a) The retrieval accuracy of the expanding EEG windows at intervals [0, t] and at intervals [t-100, t] respectively. (b) The retrieval accuracy of the expanding MEG windows. (c) Images reconstructed by EEG as the EEG window expands. When the EEG time window is greater than 200ms, the reconstructed image is reliable.
3.5 Spatial analysis
To examine the contribution of different brain regions to visual decoding, we divided the EEG electrodes from the THING-EEG data into five distinct brain regions (i.e., Frontal, Temporal, Center, Parietal, Occipital regions in Fig. 9a), and then conducted ablation experiments on retrieval task (Fig. 9b) and the reconstruction task (Fig. 9c). The results showed that using information from all brain regions is optimal, for both retrieval and generation tasks. The occipital had the highest retrieval accuracy and reconstruction performance compared to other regions. Parietal and temporal regions contain some semantic information, whereas frontal and central regions contribute the least useful information to the visual decoding.
Figure 9: EEG-guided visual retrieval and reconstruction using EEG signals from different brain regions. (a) The EEG electrodes assigned to five brain regions. (b) Top-1 and top-5 retrieval accuracy, using only the EEG channels in each leaved region and all channels. (c) Reconstructed images obtained using only the electrode channels in each individual region and all channels.
4 Related works
Visual decoding using neural signal: Decoding visual information from our brain has been a long-standing pursuit in neuroscience and computer science [30, 20]. Some progress has been made in decoding steady-state visual stimuli. However, accurately and rapidly decoding semantic information in natural images remains a challenge [42]. fMRI has been widely used to estimate semantic and shape information in visual processing within the brain [48, 19]. However, the demand for high-speed and practical applications in brain-computer interfaces calls for alternative approaches. EEG, due to its high temporal resolution and portability, emerges as a promising option [53]. Yet, the overall performance across different subjects and biological plausibility remains unresolved [1]. Furthermore, previous approaches often relied on supervised learning methods with limited image categories, overlooking the intrinsic relationship between image stimuli and brain responses [28, 41, 27].
Neural decoding for EEG signal: Previous studies have shown the efficacy of spatiotemporal modules in representing neural data [37, 25]. For example, lightweight convolutional neural networks such as EEGNet and ShallowNet  [37] have achieved considerable performance in small EEG and MEG datasets. Using contrastive learning, it has been shown that merely using convolutional neural networks and projection layers can yield satisfactory results on neural datasets [6]. More recently, Benchetrit et al proposed a B.D. encoder for MEG embedding, allowing real-time MEG-based reconstruction of visual perception [4]. Song et al. presented an EEG encoder using ShallowNet spatiotemporal convolution module with a large convolution kernel with a few parameters for EEG embedding, resulting in favorable performance on EEG-based visual decoding [46].
Limitations of previous studies: Previous EEG studies are primarily oriented toward understanding visual perception in the human brain rather than maximizing EEG decoding performance. Thus the visual decoding performance is far from optimal. Specifically, previous studies have trained linear models to (1) classify a small set of images from brain activity [15, 23], (2) to predict brain activity from the latent representations of images [8], or (3) to quantify the similarity analysis between these two patterns with representational similarity [8, 15, 14, 3]. While these studies also utilize image embeddings, their linear decoders are limited to classifying a small group of object categories or distinguishing image pairs. Moreover, several deep neural networks have been applied to maximize classification of speech [10], cognitive load [21], and images [32, 29, 2] in EEG recordings. [32] proposed a deep convolutional neural network for classifying natural images using EEG signals. Unfortunately, the experiment presented all images of the same category in a single block, probably misleading the decoder to rely on autocorrelated noise rather than the hidden informative patterns of brain activity [27]. Also, these EEG studies only classify a relatively small number of image categories.
5 Discussion and Conclusion
In this study, we developed an EEG-based image decoding and reconstruction framework. Employing self-supervised learning, our method enables the model to achieve better generalization performance in different tasks. To our knowledge, this is the first instance of EEG decoding achieving end-to-end image reconstruction with fMRI-level performance.
Technical Impact: Our technical contributions are mainly on the EEG encoder and the two-stage image generator (Fig. 2). First, we developed the ATM, a plug-and-play EEG encoder which can efficiently extract EEG/MEG features for the three visual decoding tasks. Our comprehensive experiments of the EEG encoder (Fig. 3), compared to various architectures and training methods, achieves SOTA performance across various metrics and tasks (Figs. 1b, 45). Second, the two-stage EEG guidance image generation achieves performance close to fMRI using only EEG data (Figs. 67, Table A3,A4), and this method is compatible with MEG data (Figs. 1c, 8b). By comparing EEG embeddings and image embeddings obtained through prior diffusion, we demonstrated that EEG embeddings obtained via contrastive learning alone were insufficient for generating credible images (Appendix C), suggesting that our two-stage generation strategy could effectively enhance generation performance even with limited data.
Neuroscience Insights: Our results offer insights into the relationship between brain activity and visual perception. We analyzed EEG-based visual decoding within different time windows to examine when visual information is perceived in the brain (Fig. 8). Our results revealed that visual information in EEG data is predominantly contained within the 200-400ms range (Fig. 8a), consistent with previous EEG studies [49, 16, 46]. Interestingly, the visual information in MEG data last up to 800ms, much longer than EEG (Fig. 8b), in line with the results reported by a previous MEG study [4, 46]. We also found that EEG performs better than MEG in visual tasks (See Appendix D for Table A4), which is different from other fields, such as speech decoding [10]. In addition, through ablation experiments of spatial information, we found that visual information is mainly encoded in the occipital and parietal areas (Fig. 9). Using the image reconstruction tasks, we can visualize the information encoded in the brain, providing a window for vision researchers to explore how humans perceive visual information.
Interesting Phenomena and Future Directions: Our study uncovered intriguing phenomena, sheding lights on the future direction of EEG decoding. First, there are non-negligible performance differences between cross-subject and within-subject settings. This performance gap arises from inter-subject differences in EEG signals [13, 58], likely attribute to heterogeneity in individual brain, differences in visual perception between individuals, and even shifts in noise distribution during EEG recording. To address this cross-subject challenge, it calls for more efforts on EEG encoder, such as more flexible neural network architectures or training with larger EEG dataset. Transfer learning and meta-learning are also future directions worth exploring [55, 54, 57]. On the other hand, large EEG models pre-trained on massive EEG datasets may be the ultimate solution for visual decoding tasks [9, 7]. Nevertheless, how to unify various electrode montages of different EEG datasets when pre-training large EEG models is a challenge. EEG source localization, which converts senor-level EEG signals into the standard brain source space [52, 50], might be a potential solution. Moreover, our EEG-based image reconstruction results are somewhat limited by the visual features used from CLIP. For instance, it is easy to reconstruct semantics but difficult to preserve low-level visual feature aspects. This limitation leads to generated images lacking detail and richness. Future studies could use multiple different visual features to align EEG and image data. This proposition finds support in similar methodologies applied in fMRI-based studies [40, 26], which combine high-level semantic features extracted by CLIP with low-level visual features from VAE to improve consistency.
Acknowledgements
This work is supported by the National Key R&D Program of China (2021YFF1200804),
Shenzhen Science and Technology Innovation Committee (20200925155957004,
KCXFZ20201221173400001, SGDX2020110309280100).
This work originated as a course project in the BI&AI 2023 course at SUSTech, and the author would like to thank course project members, Fan Yuanhao, Wu Zhihong, and Qin Haoyang, for their contributions to the early stages of this work.
Appendix 0.A Datasets for experiments
0.A.1 EEG dataset
We conducted our experiments on the THINGS-EEG dataset’s training set [14, 16]. This dataset includes a large EEG corpus from 10 human subjects during the visual task. The experiment employed the Rapid Serial Visual Presentation (RSVP) paradigm for orthogonal target detection tasks to ensure that participants attended to the visual stimuli. All 10 participants completed 4 equivalent experiments, resulting in 10 datasets with 16,540 training image conditions repeated 4 times, and 200 testing image conditions repeated 80 times, totaling (16,540 training image conditions $\times$ 4 repetitions) + (200 testing image conditions $\times$ 80 repetitions) = 82,160 image trials. Original data were collected using a 64-channel system at a sampling rate of 1000 Hz. After signal denoising, epoch data were downsampled to 100 Hz, selecting 17 channels covering the occipital and parietal cortex. Instead of using the raw dataset, we chose to filter it to [0.1, 100] Hz, retaining 63 channels of the original EEG data at a sampling rate of 1000 Hz. For preprocessing, we segmented the EEG data from 0 to 1000 ms after the stimulus onset into trials. Baseline correction was performed using the mean of the 200 ms pre-stimulus data. All electrodes were retained and downsampled to 250 Hz for analysis, and multivariate noise normalization was applied to the training data [17]. To improve signal-to-noise ratio, we averaged across the four EEG trials from the same image in the test set, while keeping each EEG trial in the training setting. We compared the effects of averaging across EEG trials and found it indeed improved the performance.
0.A.2 MEG dataset
To verify the versatility of ATM for embedding electrophysiological data, we tested it on MEG data modality using the THINGS-MEG dataset [18]. It includes 271-channel MEG data from 4 subjects with 12 MEG sessions. The training dataset has 1854 Concepts $\times$ 12 images $\times$ 1 repetition, and the test dataset has concepts $\times$ 1 image $\times$ 12 repetitions for 200 times. Here, we discarded 200 testing concepts from the training set to construct the same zero-shot task as with the THINGS-EEG. Each image in the THINGS-MEG was displayed for 500 ms. There was a fixed time for each image of 1000 $\pm$ 200 ms. Continuous MEG data from -100 ms to 1300 ms was segmented into trials after the stimulus onset from 0 to 1000 ms. Preprocessing was performed using a bandpass filter of [0.1, 40] Hz and baseline correction after downsampling to 200 Hz. Note that due to the small number of participants, no statistical analysis was performed on the MEG dataset. We compared our approach with advanced methods i.e. NICE [46] and B.D. [4] for classification and retrieval tasks on the MEG dataset. Similar to [46], we directly used the stimulus images to match the template, rather than other images belonging to the concept.
Appendix 0.B More Implementation Details
0.B.1 Evaluation metric implementation
Classification accuracy
As CLIP has been designed to align text and image modalities, we also leverage its text encoder for EEG classification using the text embeddings of categories. This approach utilizes CLIP’s text encoding capabilities to facilitate EEG classification. We conducted zero-shot classification tests on the THINGS-EEG dataset. We employed Top-K accuracy as a metric for performance evaluation. Specifically, we assessed performance based on the Top-k (where k=1, 5) predictions. This means a classification is considered correct if the true category is among the model’s Top-k predicted categories. We conducted tests for both within-subject and leave-one-subject-out classification accuracy, enabling a comprehensive evaluation of the model’s performance across different scenarios. Additionally, for each test instance, we extracted embeddings of N-1 unrelated samples from the test set as inputs. This means, apart from the entire test set, the model evaluated by N-Way accuracy (where N=2, 4, 10 in our experiments) on the test set. We report these results in Appendix 0.G.
Retrieval accuracy
Similar to the classification task, in the retrieval task, the objective is to retrieve the Top-K images most related to a given stimulus image via its corresponding EEG signal. This implies that by merely changing the text embeddings of image labels used during prediction, to image embeddings, we can transition the task from classification to image retrieval. The performance in retrieval tasks is superior to classification tasks as using image embeddings in training. We conducted a detailed comparison of different methods in terms of their Top-K retrieval accuracy in image retrieval tasks. Given that contrastive learning is known to be sensitive to batch size, we also compared the performance improvement of different methods under varying batch sizes (batch size=16, 1024) (Appendix 0.G).
Generation accuracy
The generation task presents more challenges than the other tasks. In this phase, EEG representations are fed into a two-stage generator. For each image condition in the test set, we generate 1 images from 10 subjects based on the corresponding EEG signal. Subsequently, image retrieval is performed for each generated image. The Top-1 and Top-5 accuracies are calculated. Since we utilized CLIP for alignment and did not extract lower-level visual features from the EEG, this metric is particularly useful. It helps in evaluating the semantic alignment between the generated images and their original counterparts.
0.B.2 Computing methods implementation
In the upstream EEG encoder part, we compared various methods. For the B.D. method  [4], we replicated the network structure as described in the original work, with the difference being in the shape of the input data, due to the original study’s focus on MEG. It is worth mentioning that we used the retention test method in the testing process, so the strategy of joint subject training is not suitable for our task. We modify its subject-wise layer as an Input layer for modeling the time dimension. Regarding the  [46], we also reproduced the EEG encoder as described in their paper, using spatiotemporal convolution modules and EEG projection modules with the same convolution parameters. To ensure fairness, we did not use the same hyperparameters as in the original paper. Instead, we chose settings that yielded excellent results upon reproduction. Across all methods, we used identical hyperparameters, apart from the network structures. These included batch size, optimizer, initial learning rate, and temperature parameters.
0.B.3 Architecture details
Table 1: Brain module configuration for use with a target latent of size 1024
Layer Input shape Output shape # parameters
Spatial attention block (N, C, T) (N, C, T) 553,078
Input layer (N, C, T) (N, C, T) 62,750
Plug-and-Play Conv module (N, C, T) (N, H1, H2) 103,680
Temporal aggregation (N, H1, H2) (N, H1*H2) 0
MLP projector (N, H1*H2) (N, 1024) 2,527,232
Total 3,246,741
Table 2: Ablation study on the ATM model’s different components for THINGS-EEG retrieval.
Module MLP AdaConv IL SA TOP-1 TOP-5
8.01 ±1.97 25.41±5.74
21.55±6.26 50.78±8.77
23.57±6.00 53.50±7.26
27.71±6.22 58.71±7.49
ATM-S 26.49±5.81 56.58±8.67
8.01 ±1.97 25.41±5.74
18.67±5.48 46.31±9.11
21.11±5.56 50.48±7.74
24.64±5.97 54.17±5.93
ATM-E 22.43±5.41 51.24±7.27
0.B.4 Training details
In the initial design of the EEG encoder module, we adopted two approaches to guide model predictions: text embedding and image embedding. Due to the differences in feature granularity, alignments favoring image embedding tend to perform better in image retrieval and classification tasks. During the training process using contrastive learning, we found that a batch size of 16 is a prudent choice across all models. A batch size of 1024 means that a sufficient number of samples are compared in a single training step, which requires the model to have a higher noise resistance capability. As observed in 5 and 6, larger batch sizes have the potential to achieve better performance and definite training efficiency gains across most methods. To improve the signal-to-noise ratio of EEG, we averaged 80 repeated data in the test set, a method similar to seeking Event-Related Potentials (ERP). To make full use of the training data volume, we did not average the 4 repetitions in the training set but instead fed all EEG data into the model for learning.
Figure 10: Test accuracy during training. (a) Training of within-subject model. (b) Training of across-subject model. We compared 5 different EEG encoding models, including EEGNetv4, B.D., NICE, ATM-S and ATM-E.
Appendix 0.C Details of EEG guidance image generation
Here, we provide a concise overview of the conditional diffusion model framework used in EEG-guided image generation, following the presentation of continuous-time diffusion models in [45, 22].
Diffusion models
Diffusion Models (DMs) engage in a generative process by transforming high-variance Gaussian noise into structured data representations. This transformation is achieved by gradually reducing noise levels across a sequence of steps. Specifically, we begin with a high-variance Gaussian noise $x_{M}\sim\mathcal{N}(0,\sigma_{\text{max}}^{2})$ and systematically denoise it through a series of steps to obtain $x_{t}\sim p(x_{t};t)$, where $\sigma_{t}<\sigma_{t+1}$ and $\sigma_{M}=\sigma_{\text{max}}$. For a well-calibrated DM, and with $\sigma_{0}=0$, the final $x_{0}$ aligns with the original data distribution.
Sampling process
The sampling in DMs is implemented by numerically simulating a Probability Flow ordinary differential equation (ODE) or a stochastic differential equation (SDE). The ODE is represented as:
$dx=-\dot{\sigma}(t)\sigma(t)\nabla_{x}\log p(x;t)dt,$ (1)
where $\nabla_{x}\log p(x;t)$ is the score function, and $\sigma(t)$ is a pre-defined schedule with its time derivative $\dot{\sigma}(t)$. The SDE variant includes a Langevin diffusion component and is expressed as:
$\begin{split}dx=&-\dot{\sigma}(t)\sigma(t)\nabla_{x}\log p(x;t)dt\\ &-\beta(t)\sigma^{2}(t)\nabla_{x}\log p(x;t)dt\\ &+\sqrt{2\beta(t)}\sigma(t)d\omega_{t},\end{split}$ (2)
where $d\omega_{t}$ is the standard Wiener process.
Training of DMs
The core of DM training is to learn a model $s_{\theta}(x;t)$ for the score function. This is typically achieved through denoising score matching (DSM), where $\epsilon_{\theta}$ is a learnable denoiser. The training process can be formulated as:
$\mathbb{E}_{(x_{0},c)\sim p_{\text{data}}(x_{0},c),(n_{t},t)\sim p(n_{t},t)}\left[\|\epsilon_{\theta}(x_{0}+n_{t};t,c)-x_{0}\|_{2}^{2}\right],$ (3)
where $n_{t}$ is Gaussian noise with variance $\sigma_{t}^{2}$, and $c$ represents a condition.
0.C.1 Stage I - EEG-Conditioned Diffusion
The initiation of the EEG-conditioned diffusion phase is paramount in our EEG-based image generation framework, leveraging the classifier-free guidance strategy alongside data pairs of CLIP embeddings and EEG embeddings $(z_{I},z_{E})$. Adapting from state-of-the-art generative techniques, our diffusion process is specifically conditioned on the EEG embedding $z_{E}$ to adeptly capture the distribution of CLIP embeddings $p(z_{I}|z_{E})$. The CLIP embedding $z_{I}$, procured during this phase, establishes the groundwork for the ensuing image generation stage. Our architecture incorporates a streamlined U-Net, labeled as $\epsilon_{\text{prior}}(z^{t}_{I},t,z_{E})$, where $z^{t}_{I}$ signifies the perturbed CLIP embedding at a given diffusion timestep $t$. The training utilizes pairs from the ImageNet database, consisting of over a million images, to fine-tune the EEG-Conditioned Diffusion model. This model is meticulously trained using the classifier-free guidance approach, effectively balancing the conditioning signal’s fidelity with the generative output’s diversity.
Classifier-free guidance method
The Classifier-Free Guidance technique is crucial in guiding the iterative refinement of a Diffusion Model (DM) under a specific EEG condition $z_{E}$. It achieves this by synchronizing the outputs of both a conditional and an unconditional model. The model’s formulation, $\epsilon^{w}_{\text{prior}}(z_{I}^{t};t,z_{E})$, is as follows:
$\epsilon^{w}_{\text{prior}}(z_{I}^{t};t,z_{E})=(1+w)\epsilon_{\text{prior}}(z_{I}^{t};t,z_{E})-w\epsilon_{\text{prior}}(z_{I}^{t};t),$ (4)
where $w\geq 0$ represents the guidance scale. This mechanism facilitates concurrent training of the conditional and unconditional models within a singular network framework, periodically substituting the EEG embedding $z_{E}$ with a null vector to promote training variability, i.e. 10% of the time. The primary objective of this method is to enhance the sample quality produced by DMs while maintaining output diversity.
0.C.2 Stage II - CLIP-Embedded Image Synthesis
In Figure 11, we compare the effects of one-stage and two-stage EEG-guided image generation. We show images generated using EEG embeddings directly (One-stage) and images generated using image embeddings obtained via prior diffusion (Two-stage). It can be seen that the two-stage EEG-guided image generation can more accurately reconstruct the semantic and low-level visual features of the original image, and the style is more realistic.
Figure 11: Comparison between one-stage and two-stage EEG guidance image generation. (a) We present the images that subjects seen (Seen), the generated images directly using EEG embeddings (One-stage), and the generated images from image embeddings obtained by the prior diffusion (Two-stage). These results indicate that the strategy of our two-stage generation can better reconstruct the seen visual stimuli. (b) We employed ATM-S to compare the generated images with the original images in a retrieval task. Our result indicates that the images generated in two stages significantly enhance the performance of the original model on the retrieval task.
In the second stage of our EEG-based image generation approach, the CLIP embedding $z_{I}$ derived from the EEG-conditioned diffusion acts as the precursor for synthesizing visual objects $I$ based on $z_{I}$. This is achieved by harnessing the synergies of advanced pre-trained models, namely SDXL and IP-Adapter [33, 59], facilitating the creation of high-caliber images.
The cornerstone of our synthesis process is the SDXL framework, acclaimed for its proficiency in text-to-image conversion. The integration of the IP-Adapter introduces dual cross-attention mechanisms, allowing the CLIP embedding $z_{I}$ to serve as a directive input and guide the denoising trajectory within the U-Net structure. The synthesis model is denoted as $\epsilon_{\text{SD}}(z_{t},t,z_{I})$, where $z_{t}$ denotes the SDXL Variational Autoencoder’s (VAE) disturbed latents.
SDXL-turbo for accelerated processing
To augment the efficiency of our framework, we additionally explore the SDXL-Turbo [36], a refined iteration of SDXL optimized for swift image synthesis. This variant proves especially beneficial in scenarios demanding quick generation of high-fidelity visuals.
IP-Adapter’s efficacy
The IP-Adapter, with its compact design, has proven to be effective in enhancing image prompt adaptability within pre-trained text-to-image models. Its compatibility with text prompts for multimodal image generation extends the versatility of our EEG-based image synthesis approach.
Appendix 0.D Performance comparison
Comparison metrics
Our study uses various metrics to evaluate how well we can recreate visual stimuli from brain data (EEG, MEG, fMRI) (Table 1 in main text). These metrics include SSIM (structural similarity index metric)[51], SwAV (SwAV-ResNet50, refer to average correlation distance)[5], and two-way identification using neural networks (AlexNet(2/5), Inception, CLIP. Here AlexNet(2/5) the 2nd and 5th feature layers of AlexNet) for both low-level and high-level image features. Here two-way identification can be seem as a two-way retrieval task described in 0.B.1. In Table 3, our results showed that on the THINGS dataset, we could achieve performance over MEG on EEG reconstruction using ATM. Table 4 shows the decoding performance of different data sets (fMRI, MEG, EEG) on visual stimulus tasks, and we even achieved the same or better performance than fMRI and MEG. Our results suggest that a suitable upstream neural representation plays a decisive role in the downstream task.
Table 3: Quantitative assessments of the reconstruction quality for EEG, MEG, and fMRI. For detailed explanations of the metrics.
Low-level High-level
Dataset $\uparrow$ SSIM $\uparrow$ AlexNet(2) $\uparrow$ AlexNet(5) $\uparrow$ Inception $\uparrow$ CLIP $\uparrow$ SwAV $\downarrow$
NSD (B.D.) [4] 0.366 0.962 0.977 0.910 0.917 0.410
NSD (Brain-Diffuser) [31] 0.356 0.942 0.962 0.872 0.915 0.423
NSD (MindEye) [39] 0.308 0.917 0.974 0.936 0.942 0.369
THINGS-MEG (B.D.) [4] 0.327 0.695 0.753 0.593 0.700 0.630
THINGS-MEG (Ours) 0.340 0.613 0.672 0.619 0.603 0.651
THINGS-EEG (Ours) 0.345 0.776 0.866 0.734 0.786 0.582
Table 4: The classification performance of various methods are discussed. Due to differences in datasets and data modalities, we have specified unified metrics to objectively assess the performance of each method.
50-way 100-way 200-way
Dataset Model top-1 top-5 top-1 top-5 top-1 top-5
GOD-Wiki (fMRI) CADA-VAE (V&T)[38] 10.02 40.37 - - - -
MVAE (V&T) [56] 10.04 39.60 - - - -
MMVAE (V&T) [43] 11.68 43.29 - - - -
MoPoE-VAE (V&T) [47] 12.90 51.78 - - - -
BraVL (V&T) [11] 13.99 53.13 - - - -
THINGS (MEG) ATM (Ours) 15.63 41.38 11.75 29.25 5.88 19.25
THINGS (EEG) BraVL [11] 14.33 40.28 - - 5.82 17.45
ATM (Ours) 17.40 39.40 11.50 28.50 7.40 20.60
Appendix 0.E Representational analysis
As depicted in Fig. 12, we showcase the representational similarity matrix and visualization in the latent space. To investigate the relationship between the representations obtained from EEG and those of images, we conducted a representational similarity matrix. We focused on subject 8, who exhibited the highest retrieval accuracy. By applying a clustering algorithm to the image embeddings corresponding to 200 images in the test set, we observed distinct within-category clustering. We generated similarity matrices based on both image and text embeddings, which were then compared with EEG representations. As shown in Fig. 12, clear within-category clustering is observable in the representational similarity matrix with image, whereas this phenomenon is not present in the representational similarity matrix with text.
Figure 12: Visualization of the representation of EEG, image and text modality. (a) Representational similarity matrix between EEG features and image/text features. (b) Visualization in the latent space of EEG/image/text by t-SNE.
Appendix 0.F Additional images results
0.F.1 Additional retrieval results
Figure 13: Additional retrieval results
0.F.2 Additional generated images
Figure 14: Additional generated results with the best alignment to original images
Figure 15: Additional generated results with the median alignment to original images
Figure 16: Additional generated results with the worst alignment to original images
0.F.3 Additional generated images for each subject
Figure 17: Part of subject 1 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Figure 18: Part of subject 2 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Figure 19: Part of subject 3 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Figure 20: Part of subject 4 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Figure 21: Part of subject 5 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Figure 22: Part of subject 6 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Figure 23: Part of subject 7 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Figure 24: Part of subject 8 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Figure 25: Part of subject 9 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Figure 26: Part of subject 10 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
Appendix 0.G Additional evaluation results
0.G.1 Accuracy for time windows
Figure 27: Accuracy for growing windows. We use an EEG time window of 100ms, sliding 100ms each time. (a) Top-1 accuracy. (b) Top-5 accuracy.
Figure 28: Accuracy for sliding windows. We use an EEG time window of 100ms, sliding 100ms each time. (a) Top-1 accuracy. (b) Top-5 accuracy.
Figure 29: Accuracy for growing windows. The MEG time window grows from 50ms to 1000ms. (a) Top-1 accuracy. (b) Top-5 accuracy.
Figure 30: Accuracy for sliding windows. We use an MEG time window of 100ms, sliding 100ms each time. (a) Top-1 accuracy. (b) Top-5 accuracy.
Table 5: Overall accuracy of zero-shot retrieval on THINGS-EEG dataset. We showed in-subject and cross-subject retrieval task performance (Ave $\pm$ Std.%) under the condition of batch size=16. We compared the 2-way, 4-way, 10-way, the Top-1 and Top-5 accuracy of 200-way from different EEG embedding methods. Our ATM outperformed all the others.
Subject dependent - train and test on one subject (batch size=16)
Methods 2-Way 4-Way 10-Way Top-1 Top-5
EEGITNet
83.23 $\pm$ 2.77
65.67 $\pm$ 3.84
43.87 $\pm$ 3.78
7.79 $\pm$ 1.27
22.12 $\pm$ 2.88
EEGConformer
89.18 $\pm$ 2.88
75.49 $\pm$ 5.30
55.32 $\pm$ 6.88
11.29 $\pm$ 3.55
33.49 $\pm$ 6.78
ShallowFBCSPNet
84.29 $\pm$ 2.56
67.13 $\pm$ 3.61
46.07 $\pm$ 4.55
8.26 $\pm$ 2.95
25.87 $\pm$ 5.03
EEGNetV4
89.03 $\pm$ 2.86
75.40 $\pm$ 5.18
56.77 $\pm$ 6.60
13.29 $\pm$ 3.99
35.50 $\pm$ 7.14
B.D.
91.14 $\pm$ 2.45
79.59 $\pm$ 4.59
62.62 $\pm$ 6.33
16.29 $\pm$ 4.35
42.16 $\pm$ 7.56
NICE
92.17 $\pm$ 2.78
82.10 $\pm$ 5.55
65.65 $\pm$ 8.03
19.40 $\pm$ 5.83
46.26 $\pm$ 10.42
MLP
87.67 $\pm$ 4.41
72.58 $\pm$ 7.62
53.52 $\pm$ 9.45
11.81 $\pm$ 4.33
32.74 $\pm$ 9.24
ATM-S (Ours)
93.89 $\pm$ 1.97
85.38 $\pm$ 3.95
71.63 $\pm$ 6.17
22.84 $\pm$ 5.87
52.22 $\pm$ 8.34
ATM-E (Ours) 94.44 $\pm$ 1.76 86.11 $\pm$ 3.39 71.89 $\pm$ 5.64 24.38 $\pm$ 6.56 54.14 $\pm$ 8.08
Subject independent - leave one subject for test (batch size=16)
Methods 2-Way 4-Way 10-Way Top-1 Top-5
EEGITNet
77.14 $\pm$ 4.14
55.29 $\pm$ 6.22
31.88 $\pm$ 6.20
2.93 $\pm$ 1.70
13.00 $\pm$ 4.45
EEGConformer
79.49 $\pm$ 3.91
59.31 $\pm$ 5.83
36.67 $\pm$ 5.35
4.11 $\pm$ 1.83
17.04 $\pm$ 4.21
ShallowFBCSPNet
75.14 $\pm$ 4.45
53.06 $\pm$ 6.70
31.23 $\pm$ 6.56
3.19 $\pm$ 2.17
13.01 $\pm$ 5.15
EEGNetV4
82.60 $\pm$ 3.17
64.28 $\pm$ 5.44
42.24 $\pm$ 6.10
6.13 $\pm$ 2.40
21.23 $\pm$ 5.19
B.D.
81.49 $\pm$ 3.52
62.35 $\pm$ 6.44
40.53 $\pm$ 7.19
6.16 $\pm$ 2.40
20.45 $\pm$ 5.40
NICE
81.85 $\pm$ 2.53
63.57 $\pm$ 4.46
41.86 $\pm$ 4.68
6.43 $\pm$ 1.46
21.39 $\pm$ 3.62
MLP
80.49 $\pm$ 2.61
62.09 $\pm$ 3.88
40.37 $\pm$ 3.71
5.67 $\pm$ 1.09
19.97 $\pm$ 3.02
ATM-S (Ours)
82.88 $\pm$ 5.18
65.03 $\pm$ 8.51
44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98
44.84 $\pm$ 7.76
7.47 $\pm$ 2.84
23.75 $\pm$ 6.86
Table 6: Overall accuracy of zero-shot Retrieval on THINGS-EEG dataset. We showed in-subject and cross-subject retrieval task performance (Ave $\pm$ Std.%) under the condition of batch size=1024. We compared the 2-way, 4-way, 10-way, the Top-1 and Top-5 accuracy of 200-way from different EEG embedding methods. Our ATM outperformed all the others.
Subject dependent - train and test on one subject (batch size=1024)
Methods 2-Way 4-Way 10-Way Top-1 Top-5
EEGITNet
76.69 $\pm$ 12.97
56.98 $\pm$ 16.31
36.35 $\pm$ 15.11
5.75 $\pm$ 3.62
18.14 $\pm$ 9.40
EEGConformer
76.17 $\pm$ 13.13
56.29 $\pm$ 16.70
34.72 $\pm$ 14.79
3.98 $\pm$ 2.80
17.10 $\pm$ 9.21
ShallowFBCSPNet
74.32 $\pm$ 12.14
53.97 $\pm$ 15.81
33.48 $\pm$ 14.35
6.10 $\pm$ 4.61
16.53 $\pm$ 9.94
EEGNetV4
92.81 $\pm$ 2.22
83.15 $\pm$ 4.20
67.81 $\pm$ 6.11
19.51 $\pm$ 5.19
48.99 $\pm$ 6.75
B.D.
78.42 $\pm$ 8.81
58.24 $\pm$ 12.13
37.97 $\pm$ 11.38
5.88 $\pm$ 3.49
18.61 $\pm$ 7.81
NICE
92.73 $\pm$ 2.75
83.26 $\pm$ 5.47
67.96 $\pm$ 8.31
19.32 $\pm$ 5.33
49.26 $\pm$ 9.69
MLP
83.09 $\pm$ 2.54
66.70 $\pm$ 4.16
45.43 $\pm$ 4.58
7.23 $\pm$ 1.66
25.14 $\pm$ 3.66
ATM-S (Ours) 94.60 $\pm$ 1.93 86.88 $\pm$ 4.20 73.89 $\pm$ 5.93 26.09 $\pm$ 6.96 58.07 $\pm$ 8.16
ATM-E (Ours)
92.99 $\pm$ 2.20
83.81 $\pm$ 4.46
68.87 $\pm$ 7.27
22.40 $\pm$ 6.62
50.59 $\pm$ 9.59
Subject independent - leave one subject for test (batch size=1024)
Methods 2-Way 4-Way 10-Way Top-1 Top-5
EEGITNet
77.47 $\pm$ 3.75
55.75 $\pm$ 6.01
33.60 $\pm$ 6.16
3.80 $\pm$ 2.01
14.02 $\pm$ 3.89
EEGConformer
68.02 $\pm$ 7.75
44.24 $\pm$ 9.25
23.84 $\pm$ 7.65
1.66 $\pm$ 1.37
8.62 $\pm$ 4.20
ShallowFBCSPNet
76.40 $\pm$ 4.64
54.81 $\pm$ 7.09
32.37 $\pm$ 7.28
2.49 $\pm$ 1.57
13.24 $\pm$ 5.98
EEGNetV4
82.60 $\pm$ 3.17
64.28 $\pm$ 5.44
42.24 $\pm$ 6.10
6.13 $\pm$ 2.40
21.23 $\pm$ 5.19
B.D.
85.30 $\pm$ 6.24
69.87 $\pm$ 10.09
50.82 $\pm$ 11.74
10.57 $\pm$ 5.26
29.74 $\pm$ 11.28
NICE
83.75 $\pm$ 3.21
65.82 $\pm$ 5.93
44.41 $\pm$ 6.12
7.04 $\pm$ 2.83
23.55 $\pm$ 4.86
MLP
80.49 $\pm$ 2.61
62.09 $\pm$ 3.88
40.37 $\pm$ 3.71
5.67 $\pm$ 1.09
19.97 $\pm$ 3.02
ATM-S (Ours)
87.36 $\pm$ 3.97
72.80 $\pm$ 7.02
53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51
53.15 $\pm$ 6.52
11.12 $\pm$ 3.26
32.61 $\pm$ 6.78
Table 7: Accuracy of zero-shot Retrieval on THINGS-EEG for odd-numbered subjects(batch size=16).
Subject dependent - train and test on one subject (batch size=16)
Method Subject 1 Subject 3 Subject 5 Subject 7 Subject 9
top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5
EEGITNet 8.20 19.45 6.95 24.25 6.30 19.50 8.95 19.00 6.10 21.60
Conformer 4.75 21.30 9.90 34.15 7.55 28.05 12.25 35.35 13.95 34.10
ShallowFBCSPNet 7.00 20.90 8.45 28.45 7.75 20.80 6.35 27.60 2.80 18.00
EEGNetV4 10.85 31.10 17.65 41.20 6.75 21.80 14.15 39.05 10.40 30.50
B.D. 13.05 36.00 14.45 37.85 10.55 32.00 18.65 46.80 14.60 37.55
NICE 13.25 35.75 23.15 51.40 12.40 33.45 19.45 50.20 17.55 41.70
MLP 13.00 38.30 13.90 39.60 8.00 24.85 13.00 39.15 12.65 36.20
ATM-S (Ours) 19.10 49.05 22.95 56.30 16.10 39.95 25.75 55.40 21.50 48.15
ATM-E (Ours) 20.75 47.35 25.10 56.45 18.50 44.35 24.70 54.10 23.40 52.40
Subject independent - leave one subject for test (batch size=16)
Method Subject 1 Subject 3 Subject 5 Subject 7 Subject 9
top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5
EEGITNet 3.40 13.75 1.05 10.80 1.45 7.55 4.30 13.65 0.55 5.95
Conformer 5.45 20.10 2.00 12.75 2.90 12.95 5.50 16.45 1.05 9.90
ShallowFBCSPNet 4.25 13.30 1.45 7.65 0.90 7.40 3.75 13.75 0.65 5.95
EEGNetV4 7.20 22.40 4.60 17.80 3.00 15.75 3.80 18.95 3.40 12.70
B.D. 6.55 21.45 5.45 17.00 3.10 14.75 6.45 24.00 2.60 12.75
NICE 4.80 20.05 5.80 18.10 5.85 16.95 5.75 21.20 4.95 17.65
MLP 3.65 16.20 5.95 20.95 5.90 19.20 4.40 14.25 7.45 23.85
ATM-S (Ours) 8.30 25.50 11.40 31.30 5.95 21.25 4.80 19.45 4.05 13.30
ATM-E (Ours) 7.80 22.95 13.75 35.70 4.20 14.30 6.40 25.95 4.50 14.90
Table 8: Accuracy of zero-shot Retrieval on THINGS-EEG for even-numbered subjects(batch size=16).
Subject dependent - train and test on one subject (batch size=16)
Method Subject 2 Subject 4 Subject 6 Subject 8 Subject 10 Overall Avg
top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5
EEGITNet 6.55 19.05 7.90 23.10 8.10 23.10 9.80 27.35 9.05 24.85 8.08 23.49
Conformer 8.10 25.80 13.30 37.10 12.75 36.65 15.75 45.10 14.55 37.35 12.89 36.40
ShallowFBCSPNet 6.15 23.50 10.85 29.10 9.05 26.95 13.15 34.90 11.00 28.55 10.04 28.60
EEGNetV4 8.80 28.20 14.30 37.45 15.40 42.45 19.40 42.40 15.15 40.85 14.61 38.27
B.D. 13.35 38.00 17.45 45.00 14.25 40.35 24.50 54.65 22.10 53.45 18.33 46.29
NICE 11.45 31.05 20.75 48.95 22.25 50.35 29.75 62.40 24.00 57.40 21.64 50.03
MLP 10.85 33.10 15.20 37.70 12.15 35.05 18.65 47.55 18.30 45.55 15.03 39.79
ATM-S (Ours) 15.10 41.00 25.75 55.75 20.05 50.10 34.90 67.50 27.15 58.95 24.59 54.66
ATM-E (Ours) 15.25 44.90 25.75 55.75 23.00 53.95 39.40 70.25 29.60 63.75 26.26 57.36
Subject independent - leave one subject for test (batch size=16)
Method Subject 2 Subject 4 Subject 6 Subject 8 Subject 10 Overall Avg
top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5
EEGITNet 2.65 11.40 3.25 15.35 8.10 23.10 9.80 27.35 9.05 24.85 6.57 20.41
Conformer 4.65 18.70 4.25 19.60 12.75 36.65 15.75 45.10 14.55 37.35 10.39 31.48
ShallowFBCSPNet 2.75 13.70 5.55 15.45 9.05 26.95 13.15 34.90 11.00 28.55 8.30 23.91
EEGNetV4 7.80 24.30 7.35 23.90 15.40 42.45 19.40 42.40 15.15 40.85 13.02 34.78
B.D. 5.00 16.40 6.65 22.35 14.25 40.35 24.50 54.65 22.10 53.45 14.50 37.44
NICE 5.55 21.40 8.20 24.90 22.25 50.35 29.75 62.40 24.00 57.40 17.95 43.29
MLP 5.50 20.65 5.45 18.35 12.15 35.05 18.65 47.55 18.30 45.55 12.01 33.43
ATM-S (Ours) 7.25 23.65 14.55 39.80 20.05 50.10 34.90 67.50 27.15 58.95 20.78 47.90
ATM-E (Ours) 8.65 27.60 14.55 39.80 23.00 53.95 39.40 70.25 29.60 63.75 21.62 47.93
Table 9: Accuracy of zero-shot Retrieval on THINGS-EEG for odd-numbered subjects(batch size=1024).
Subject dependent - train and test on one subject (batch size=1024)
Method Subject 1 Subject 3 Subject 5 Subject 7 Subject 9
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
EEGITNet 0.50 3.45 2.40 7.75 5.35 18.25 8.90 24.90 5.50 22.45
Conformer 0.50 3.20 0.50 6.90 3.40 17.50 7.10 23.80 6.15 22.75
ShallowFBCSPNet 0.65 3.35 2.30 6.20 3.35 14.55 7.35 23.15 6.80 18.05
EEGNetV4 11.45 41.45 20.85 48.70 13.75 42.00 19.05 50.30 15.50 40.60
B.D. 6.70 23.25 6.70 19.75 1.70 8.10 5.85 20.55 6.80 21.05
NICE 17.65 50.25 22.05 51.20 9.80 27.65 17.15 49.60 19.60 52.55
MLP 8.40 24.40 5.60 22.65 6.00 25.15 8.35 24.35 5.45 22.25
ATM-S (Ours) 25.60 60.40 25.00 62.35 12.90 43.00 30.50 61.50 24.35 51.50
ATM-E (Ours) 18.85 44.10 24.30 56.10 11.95 32.65 24.85 53.75 20.05 47.75
Subject independent - leave one subject for test (batch size=1024)
Method Subject 1 Subject 3 Subject 5 Subject 7 Subject 9
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
EEGITNet 4.45 13.95 1.50 12.05 1.00 8.60 4.50 13.35 1.00 8.95
Conformer 4.45 17.20 0.70 4.75 0.50 5.30 2.40 10.95 0.50 3.35
ShallowFBCSPNet 2.00 11.25 1.70 9.25 1.00 4.85 3.80 17.30 1.00 5.75
EEGNetV4 9.45 24.00 8.20 21.90 5.20 18.75 5.45 21.80 5.65 16.40
B.D. 5.30 16.60 13.45 35.30 8.15 21.30 15.85 37.75 8.10 27.10
NICE 4.80 20.05 6.75 20.70 4.15 18.55 5.55 24.45 5.10 17.20
MLP 4.45 12.30 6.75 24.60 4.90 19.55 4.85 17.60 4.45 19.80
ATM-S (Ours) 10.45 26.75 11.85 33.80 6.95 23.85 16.05 43.50 4.85 22.70
ATM-E (Ours) 10.45 26.75 13.45 35.30 8.15 24.75 15.85 39.05 7.55 30.65
Table 10: Accuracy of zero-shot Retrieval on THINGS-EEG for even-numbered subjects(batch size=1024).
Subject dependent - train and test on one subject (batch size=1024)
Method Subject 2 Subject 4 Subject 6 Subject 8 Subject 10 Overall Avg
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
EEGITNet 0.75 5.55 6.00 20.35 7.60 20.80 10.55 28.00 9.90 29.95 7.16 20.93
Conformer 0.50 3.70 3.30 18.10 4.35 21.95 7.50 25.00 6.50 28.15 4.43 19.38
ShallowFBCSPNet 0.75 4.75 5.25 16.35 7.80 16.95 13.45 32.20 13.25 29.70 8.10 20.19
EEGNetV4 17.30 43.60 27.20 58.75 20.55 53.25 22.95 56.45 26.50 54.85 22.90 53.38
B.D. 1.55 6.90 3.65 15.30 3.55 13.05 9.65 27.00 12.70 31.15 6.22 18.68
NICE 14.35 40.00 20.00 49.90 18.35 49.75 28.05 62.65 26.25 59.10 21.40 52.28
MLP 6.40 19.45 7.35 27.50 6.75 25.40 7.00 27.20 10.95 33.05 7.69 26.52
ATM-S (Ours) 22.00 54.50 31.35 60.90 21.30 51.05 38.80 72.00 29.05 63.50 28.50 60.39
ATM-E (Ours) 17.00 44.85 23.00 51.25 19.40 47.55 35.65 66.75 29.00 61.15 24.81 54.31
Subject independent - leave one subject for test (batch size=1024)
Method Subject 2 Subject 4 Subject 6 Subject 8 Subject 10 Overall Avg
EEGITNet 4.70 15.20 4.80 14.40 5.50 17.45 3.70 14.25 6.90 22.00 5.12 16.66
Conformer 1.40 9.70 1.15 8.20 1.15 9.70 0.80 5.15 3.55 11.95 1.61 8.94
ShallowFBCSPNet 2.50 12.15 2.60 17.20 2.15 16.95 1.80 13.15 6.30 24.55 3.07 16.80
EEGNetV4 8.25 28.15 9.45 26.05 6.40 24.60 9.20 23.60 10.15 36.95 8.69 27.87
B.D. 4.45 23.20 4.30 12.90 11.15 33.70 18.05 46.30 16.90 43.30 10.97 31.88
NICE 4.90 22.80 9.65 29.70 8.35 24.35 7.90 25.00 13.25 32.75 8.81 26.92
MLP 6.50 21.55 9.50 25.15 3.60 16.15 7.25 29.25 12.80 34.35 7.93 25.29
ATM-S (Ours) 7.10 24.75 14.65 39.40 11.10 35.80 14.95 40.25 20.45 46.50 13.65 37.34
ATM-E (Ours) 10.45 27.75 14.65 35.25 10.35 35.20 13.95 38.60 15.85 42.85 13.05 35.93

### v3

@@
-Visual Decoding and Reconstruction via EEG Embeddings with Guided Diffusion
+Shenzhen, China
 Abstract
@@
 This stage mainly focuses on converting CLIP embeddings into corresponding images.
-Further details are provided in Appendix C.
+Further details are provided in Appendix 0.C.
 2.4 Loss Function
@@
 (c) Images showing the top-5 accuracy in EEG-image retrieval tasks.
-See Appendix E for additional images results.
+See Appendix 0.F for additional images results.
 (d) Representation similarity analysis.
 Clustering of image features and EEG features using a k-means algorithm with k=5 is performed, followed by the calculation of similarity between clusters.
-See Appendix D for additional Representational similarity results.
+See Appendix 0.E for additional Representational similarity results.
 3 Experiments
@@
 For fairness, all models’ hyperparameters were kept consistent.
-In our study, we compared the performance of different encoders on the within-subject test set and cross-subject (leave-one-subject-out) test set (see Appendix F).
+In our study, we compared the performance of different encoders on the within-subject test set and cross-subject (leave-one-subject-out) test set (see Appendix 0.G).
 3.2 EEG Decoding performance
@@
 4a).
-More details of the EEG-based image classification are in Appendix B.
+More details of the EEG-based image classification are in Appendix 0.B.
 Fig.
-4b presents the accuracy of EEG classification for each subject, and Fig.
-4c shows our method outperforms others.
+4b presents the average accuracy across different methods in the subjects, and shows our method outperforms others.
 Here, we test the effectiveness of ATM-extracted EEG embeddings in the image retrieval task.
@@
 Fig.
-5b shows the Top-5 retrieved images corresponding to the real visual stimuli seen by subjects.
-To be a fair comparison, here we presented the images that the previous method failed to retrieve in Top-1 for testing [46].
+5b shows the average results for all subjects in the in-subject test.
+We take the highest test accuracy in the training process as the statistical result.
+See the Appendix 0.G for more detailed averages of test accuracy in subjects.
+Fig.
+5c shows the Top-5 retrieved images corresponding to the real visual stimuli seen by subjects.
 Compared with the previous model, the Top-1 accuracy of our model is significantly improved, and the Top-5 images all maintain a high degree of similarity with the original images.
-Fig.
-5c shows our method outperforms others.
+We calculated the similarity of the EEG to the semantics and images.
+The similarity between the image and EEG is shown in Fig.
+5d.
+We performed a representational similarity analysis, and we could observe different degrees of intra-class aggregation.
+This means that the EEG representation obtained is closer to the representation of the corresponding image itself than to the semantics.
 Figure 6: EEG guidance image generation.
@@
 (c) Comparison between the original image and the image generated using the corresponding EEG data.
-(see Appendix B & C for details).
+(see Appendix 0.C for details).
 (d) The similarity between visual objects and target EEG embeddings as the guidance scale changes, and the diversity of visual objects as the guidance scale changes.
-See Appendix E for additional results.
+See Appendix 0.F for additional results.
 Ablation study on ATM
 We systematically deconstructed and analyzed each layer of our EEG encoder module.
-We conducted an ablation study for each component in ATM (i.e., the MLP projector, the adaptive convolution module, the subject-wise linear, and the spatial attention block).
+We conducted an ablation study for each component in ATM (i.e., the MLP projector, the adaptive convolution module, the Input layer, and the spatial attention block).
 Notably, ATM is a flexible, plug-and-play module, which can be easily replaced with any form of convolutional component or even omitted.
-Appendix B.3 showed that the adaptive convolutional components significantly enhanced encoder performance, while the subject-wise linear layer reduced the model’s training consumption and substantially improved the existing components’ capabilities.
-Surprisingly, our spatial attention module based on sine-cosine positional encoding improved encoder performance only when used alone, but the benefits were not as significant when combined with the subject-wise linear layer.
+Appendix 0.B.3 showed that the adaptive convolutional components significantly enhanced encoder performance, while the Input layer reduced the model’s training consumption and substantially improved the existing components’ capabilities.
+Surprisingly, our spatial attention module based on sine-cosine positional encoding improved encoder performance only when used alone, but the benefits were not as significant when combined with the Input layer.
 3.3 Image Generation performance
@@
 We show the images subjects seen and the generated images by our two-stage image generator.
-See Appendix E for additional results.
+See Appendix 0.F for additional results.
 Fig.
@@
 6d).
-Surprisingly, the retrieval accuracy using generated images is even higher than that using EEG embeddings, implying the benefit of our two-stage visual reconstruction for image retrieval (See Appendix C for Figure A2).
-We also report the decoding and reconstruction performance for EEG, MEG, and fMRI across various metrics in the Appendix D.
+Surprisingly, the retrieval accuracy using generated images is even higher than that using EEG embeddings, implying the benefit of our two-stage visual reconstruction for image retrieval (See Appendix 0.C for Figure 11).
+We also report the decoding and reconstruction performance for EEG, MEG, and fMRI across various metrics in the Appendix 0.D.
 We visualize the best, medium and worst generated images in Fig.
@@
 Second, the two-stage EEG guidance image generation achieves performance close to fMRI using only EEG data (Figs.
-67, Table A3,A4), and this method is compatible with MEG data (Figs.
+67, Tab.
+3, 4), and this method is compatible with MEG data (Figs.
 1c, 8b).
-By comparing EEG embeddings and image embeddings obtained through prior diffusion, we demonstrated that EEG embeddings obtained via contrastive learning alone were insufficient for generating credible images (Appendix C), suggesting that our two-stage generation strategy could effectively enhance generation performance even with limited data.
+By comparing EEG embeddings and image embeddings obtained through prior diffusion, we demonstrated that EEG embeddings obtained via contrastive learning alone were insufficient for generating credible images (Appendix 0.C), suggesting that our two-stage generation strategy could effectively enhance generation performance even with limited data.
 Neuroscience Insights: Our results offer insights into the relationship between brain activity and visual perception.
@@
 8b), in line with the results reported by a previous MEG study [4, 46].
-We also found that EEG performs better than MEG in visual tasks (See Appendix D for Table A4), which is different from other fields, such as speech decoding [10].
+We also found that EEG performs better than MEG in visual tasks (See Appendix 0.D for Tab.
+4), which is different from other fields, such as speech decoding [10].
 In addition, through ablation experiments of spatial information, we found that visual information is mainly encoded in the occipital and parietal areas (Fig.
@@
 KCXFZ20201221173400001, SGDX2020110309280100).
-This work originated as a course project in the BI&AI 2023 course at SUSTech, and the author would like to thank course project members, Fan Yuanhao, Wu Zhihong, and Qin Haoyang, for their contributions to the early stages of this work.
+This work originated as a course project in the BI&AI 2023 course at SUSTech, and the author would like to thank course project members, Yuanhao Fan and Zhihong Wu, for their contributions to the early stages of this work.
+Thank Haoyang Qin for his time and advice.
 Appendix 0.A Datasets for experiments
@@
 A batch size of 1024 means that a sufficient number of samples are compared in a single training step, which requires the model to have a higher noise resistance capability.
-As observed in 5 and 6, larger batch sizes have the potential to achieve better performance and definite training efficiency gains across most methods.
+As observed in Tab.
+5 and Tab.
+6, larger batch sizes have the potential to achieve better performance and definite training efficiency gains across most methods.
 To improve the signal-to-noise ratio of EEG, we averaged 80 repeated data in the test set, a method similar to seeking Event-Related Potentials (ERP).
@@
 Here AlexNet(2/5) the 2nd and 5th feature layers of AlexNet) for both low-level and high-level image features.
-Here two-way identification can be seem as a two-way retrieval task described in 0.B.1.
-In Table 3, our results showed that on the THINGS dataset, we could achieve performance over MEG on EEG reconstruction using ATM.
-Table 4 shows the decoding performance of different data sets (fMRI, MEG, EEG) on visual stimulus tasks, and we even achieved the same or better performance than fMRI and MEG.
+Here two-way identification can be seem as a two-way retrieval task described in Appendix  0.B.1.
+In Tab.
+3, our results showed that on the THINGS dataset, we could achieve performance over MEG on EEG reconstruction using ATM.
+Tab.
+4 shows the decoding performance of different data sets (fMRI, MEG, EEG) on visual stimulus tasks, and we even achieved the same or better performance than fMRI and MEG.
 Our results suggest that a suitable upstream neural representation plays a decisive role in the downstream task.

### v4

@@
 It employs a prior diffusion model that is conditioned on the EEG embeddings to generate the image embeddings, followed an enhanced Stable Diffusion module to generate images with EEG priors.
-4
 Figure 3: The structure of EEG encoder ATM.
@@
 6d).
-Surprisingly, the retrieval accuracy using generated images is even higher than that using EEG embeddings, implying the benefit of our two-stage visual reconstruction for image retrieval (See Appendix 0.C for Figure 11).
+Surprisingly, the retrieval accuracy using generated images is even higher than that using EEG embeddings, implying the benefit of our two-stage visual reconstruction for image retrieval (See Appendix 0.C for Fig.11).
 We also report the decoding and reconstruction performance for EEG, MEG, and fMRI across various metrics in the Appendix 0.D.
@@
 0.C.2 Stage II - CLIP-Embedded Image Synthesis
-In Figure 11, we compare the effects of one-stage and two-stage EEG-guided image generation.
+In Fig.
+11, we compare the effects of one-stage and two-stage EEG-guided image generation.
 We show images generated using EEG embeddings directly (One-stage) and images generated using image embeddings obtained via prior diffusion (Two-stage).
@@
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet
-83.23 $\pm$ 2.77
-65.67 $\pm$ 3.84
-43.87 $\pm$ 3.78
-7.79 $\pm$ 1.27
-22.12 $\pm$ 2.88
-EEGConformer
-89.18 $\pm$ 2.88
-75.49 $\pm$ 5.30
-55.32 $\pm$ 6.88
-11.29 $\pm$ 3.55
-33.49 $\pm$ 6.78
-ShallowFBCSPNet
-84.29 $\pm$ 2.56
-67.13 $\pm$ 3.61
-46.07 $\pm$ 4.55
-8.26 $\pm$ 2.95
-25.87 $\pm$ 5.03
-EEGNetV4
-89.03 $\pm$ 2.86
-75.40 $\pm$ 5.18
-56.77 $\pm$ 6.60
-13.29 $\pm$ 3.99
-35.50 $\pm$ 7.14
-B.D.
-91.14 $\pm$ 2.45
-79.59 $\pm$ 4.59
-62.62 $\pm$ 6.33
-16.29 $\pm$ 4.35
-42.16 $\pm$ 7.56
-NICE
-92.17 $\pm$ 2.78
-82.10 $\pm$ 5.55
-65.65 $\pm$ 8.03
-19.40 $\pm$ 5.83
-46.26 $\pm$ 10.42
-MLP
-87.67 $\pm$ 4.41
-72.58 $\pm$ 7.62
-53.52 $\pm$ 9.45
-11.81 $\pm$ 4.33
-32.74 $\pm$ 9.24
-ATM-S (Ours)
-93.89 $\pm$ 1.97
-85.38 $\pm$ 3.95
-71.63 $\pm$ 6.17
-22.84 $\pm$ 5.87
-52.22 $\pm$ 8.34
+EEGITNet 83.23 $\pm$ 2.77 65.67 $\pm$ 3.84 43.87 $\pm$ 3.78 7.79 $\pm$ 1.27 22.12 $\pm$ 2.88
+EEGConformer 89.18 $\pm$ 2.88 75.49 $\pm$ 5.30 55.32 $\pm$ 6.88 11.29 $\pm$ 3.55 33.49 $\pm$ 6.78
+ShallowFBCSPNet 84.29 $\pm$ 2.56 67.13 $\pm$ 3.61 46.07 $\pm$ 4.55 8.26 $\pm$ 2.95 25.87 $\pm$ 5.03
+EEGNetV4 89.03 $\pm$ 2.86 75.40 $\pm$ 5.18 56.77 $\pm$ 6.60 13.29 $\pm$ 3.99 35.50 $\pm$ 7.14
+B.D. 91.14 $\pm$ 2.45 79.59 $\pm$ 4.59 62.62 $\pm$ 6.33 16.29 $\pm$ 4.35 42.16 $\pm$ 7.56
+NICE 92.17 $\pm$ 2.78 82.10 $\pm$ 5.55 65.65 $\pm$ 8.03 19.40 $\pm$ 5.83 46.26 $\pm$ 10.42
+MLP 87.67 $\pm$ 4.41 72.58 $\pm$ 7.62 53.52 $\pm$ 9.45 11.81 $\pm$ 4.33 32.74 $\pm$ 9.24
+ATM-S (Ours) 93.89 $\pm$ 1.97 85.38 $\pm$ 3.95 71.63 $\pm$ 6.17 22.84 $\pm$ 5.87 52.22 $\pm$ 8.34
 ATM-E (Ours) 94.44 $\pm$ 1.76 86.11 $\pm$ 3.39 71.89 $\pm$ 5.64 24.38 $\pm$ 6.56 54.14 $\pm$ 8.08
@@
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet
-77.14 $\pm$ 4.14
-55.29 $\pm$ 6.22
-31.88 $\pm$ 6.20
-2.93 $\pm$ 1.70
-13.00 $\pm$ 4.45
-EEGConformer
-79.49 $\pm$ 3.91
-59.31 $\pm$ 5.83
-36.67 $\pm$ 5.35
-4.11 $\pm$ 1.83
-17.04 $\pm$ 4.21
-ShallowFBCSPNet
-75.14 $\pm$ 4.45
-53.06 $\pm$ 6.70
-31.23 $\pm$ 6.56
-3.19 $\pm$ 2.17
-13.01 $\pm$ 5.15
-EEGNetV4
-82.60 $\pm$ 3.17
-64.28 $\pm$ 5.44
-42.24 $\pm$ 6.10
-6.13 $\pm$ 2.40
-21.23 $\pm$ 5.19
-B.D.
-81.49 $\pm$ 3.52
-62.35 $\pm$ 6.44
-40.53 $\pm$ 7.19
-6.16 $\pm$ 2.40
-20.45 $\pm$ 5.40
-NICE
-81.85 $\pm$ 2.53
-63.57 $\pm$ 4.46
-41.86 $\pm$ 4.68
-6.43 $\pm$ 1.46
-21.39 $\pm$ 3.62
-MLP
-80.49 $\pm$ 2.61
-62.09 $\pm$ 3.88
-40.37 $\pm$ 3.71
-5.67 $\pm$ 1.09
-19.97 $\pm$ 3.02
-ATM-S (Ours)
-82.88 $\pm$ 5.18
-65.03 $\pm$ 8.51
-44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
-ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98
-44.84 $\pm$ 7.76
-7.47 $\pm$ 2.84
-23.75 $\pm$ 6.86
+EEGITNet 77.14 $\pm$ 4.14 55.29 $\pm$ 6.22 31.88 $\pm$ 6.20 2.93 $\pm$ 1.70 13.00 $\pm$ 4.45
+EEGConformer 79.49 $\pm$ 3.91 59.31 $\pm$ 5.83 36.67 $\pm$ 5.35 4.11 $\pm$ 1.83 17.04 $\pm$ 4.21
+ShallowFBCSPNet 75.14 $\pm$ 4.45 53.06 $\pm$ 6.70 31.23 $\pm$ 6.56 3.19 $\pm$ 2.17 13.01 $\pm$ 5.15
+EEGNetV4 82.60 $\pm$ 3.17 64.28 $\pm$ 5.44 42.24 $\pm$ 6.10 6.13 $\pm$ 2.40 21.23 $\pm$ 5.19
+B.D. 81.49 $\pm$ 3.52 62.35 $\pm$ 6.44 40.53 $\pm$ 7.19 6.16 $\pm$ 2.40 20.45 $\pm$ 5.40
+NICE 81.85 $\pm$ 2.53 63.57 $\pm$ 4.46 41.86 $\pm$ 4.68 6.43 $\pm$ 1.46 21.39 $\pm$ 3.62
+MLP 80.49 $\pm$ 2.61 62.09 $\pm$ 3.88 40.37 $\pm$ 3.71 5.67 $\pm$ 1.09 19.97 $\pm$ 3.02
+ATM-S (Ours) 82.88 $\pm$ 5.18 65.03 $\pm$ 8.51 44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
+ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98 44.84 $\pm$ 7.76 7.47 $\pm$ 2.84 23.75 $\pm$ 6.86
 Table 6: Overall accuracy of zero-shot Retrieval on THINGS-EEG dataset.
@@
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet
-76.69 $\pm$ 12.97
-56.98 $\pm$ 16.31
-36.35 $\pm$ 15.11
-5.75 $\pm$ 3.62
-18.14 $\pm$ 9.40
-EEGConformer
-76.17 $\pm$ 13.13
-56.29 $\pm$ 16.70
-34.72 $\pm$ 14.79
-3.98 $\pm$ 2.80
-17.10 $\pm$ 9.21
-ShallowFBCSPNet
-74.32 $\pm$ 12.14
-53.97 $\pm$ 15.81
-33.48 $\pm$ 14.35
-6.10 $\pm$ 4.61
-16.53 $\pm$ 9.94
-EEGNetV4
-92.81 $\pm$ 2.22
-83.15 $\pm$ 4.20
-67.81 $\pm$ 6.11
-19.51 $\pm$ 5.19
-48.99 $\pm$ 6.75
-B.D.
-78.42 $\pm$ 8.81
-58.24 $\pm$ 12.13
-37.97 $\pm$ 11.38
-5.88 $\pm$ 3.49
-18.61 $\pm$ 7.81
-NICE
-92.73 $\pm$ 2.75
-83.26 $\pm$ 5.47
-67.96 $\pm$ 8.31
-19.32 $\pm$ 5.33
-49.26 $\pm$ 9.69
-MLP
-83.09 $\pm$ 2.54
-66.70 $\pm$ 4.16
-45.43 $\pm$ 4.58
-7.23 $\pm$ 1.66
-25.14 $\pm$ 3.66
+EEGITNet 76.69 $\pm$ 12.97 56.98 $\pm$ 16.31 36.35 $\pm$ 15.11 5.75 $\pm$ 3.62 18.14 $\pm$ 9.40
+EEGConformer 76.17 $\pm$ 13.13 56.29 $\pm$ 16.70 34.72 $\pm$ 14.79 3.98 $\pm$ 2.80 17.10 $\pm$ 9.21
+ShallowFBCSPNet 74.32 $\pm$ 12.14 53.97 $\pm$ 15.81 33.48 $\pm$ 14.35 6.10 $\pm$ 4.61 16.53 $\pm$ 9.94
+EEGNetV4 92.81 $\pm$ 2.22 83.15 $\pm$ 4.20 67.81 $\pm$ 6.11 19.51 $\pm$ 5.19 48.99 $\pm$ 6.75
+B.D. 78.42 $\pm$ 8.81 58.24 $\pm$ 12.13 37.97 $\pm$ 11.38 5.88 $\pm$ 3.49 18.61 $\pm$ 7.81
+NICE 92.73 $\pm$ 2.75 83.26 $\pm$ 5.47 67.96 $\pm$ 8.31 19.32 $\pm$ 5.33 49.26 $\pm$ 9.69
+MLP 83.09 $\pm$ 2.54 66.70 $\pm$ 4.16 45.43 $\pm$ 4.58 7.23 $\pm$ 1.66 25.14 $\pm$ 3.66
 ATM-S (Ours) 94.60 $\pm$ 1.93 86.88 $\pm$ 4.20 73.89 $\pm$ 5.93 26.09 $\pm$ 6.96 58.07 $\pm$ 8.16
-ATM-E (Ours)
-92.99 $\pm$ 2.20
-83.81 $\pm$ 4.46
-68.87 $\pm$ 7.27
-22.40 $\pm$ 6.62
-50.59 $\pm$ 9.59
+ATM-E (Ours) 92.99 $\pm$ 2.20 83.81 $\pm$ 4.46 68.87 $\pm$ 7.27 22.40 $\pm$ 6.62 50.59 $\pm$ 9.59
 Subject independent - leave one subject for test (batch size=1024)
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet
-77.47 $\pm$ 3.75
-55.75 $\pm$ 6.01
-33.60 $\pm$ 6.16
-3.80 $\pm$ 2.01
-14.02 $\pm$ 3.89
-EEGConformer
-68.02 $\pm$ 7.75
-44.24 $\pm$ 9.25
-23.84 $\pm$ 7.65
-1.66 $\pm$ 1.37
-8.62 $\pm$ 4.20
-ShallowFBCSPNet
-76.40 $\pm$ 4.64
-54.81 $\pm$ 7.09
-32.37 $\pm$ 7.28
-2.49 $\pm$ 1.57
-13.24 $\pm$ 5.98
-EEGNetV4
-82.60 $\pm$ 3.17
-64.28 $\pm$ 5.44
-42.24 $\pm$ 6.10
-6.13 $\pm$ 2.40
-21.23 $\pm$ 5.19
-B.D.
-85.30 $\pm$ 6.24
-69.87 $\pm$ 10.09
-50.82 $\pm$ 11.74
-10.57 $\pm$ 5.26
-29.74 $\pm$ 11.28
-NICE
-83.75 $\pm$ 3.21
-65.82 $\pm$ 5.93
-44.41 $\pm$ 6.12
-7.04 $\pm$ 2.83
-23.55 $\pm$ 4.86
-MLP
-80.49 $\pm$ 2.61
-62.09 $\pm$ 3.88
-40.37 $\pm$ 3.71
-5.67 $\pm$ 1.09
-19.97 $\pm$ 3.02
-ATM-S (Ours)
-87.36 $\pm$ 3.97
-72.80 $\pm$ 7.02
-53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
-ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51
-53.15 $\pm$ 6.52
-11.12 $\pm$ 3.26
-32.61 $\pm$ 6.78
+EEGITNet 77.47 $\pm$ 3.75 55.75 $\pm$ 6.01 33.60 $\pm$ 6.16 3.80 $\pm$ 2.01 14.02 $\pm$ 3.89
+EEGConformer 68.02 $\pm$ 7.75 44.24 $\pm$ 9.25 23.84 $\pm$ 7.65 1.66 $\pm$ 1.37 8.62 $\pm$ 4.20
+ShallowFBCSPNet 76.40 $\pm$ 4.64 54.81 $\pm$ 7.09 32.37 $\pm$ 7.28 2.49 $\pm$ 1.57 13.24 $\pm$ 5.98
+EEGNetV4 82.60 $\pm$ 3.17 64.28 $\pm$ 5.44 42.24 $\pm$ 6.10 6.13 $\pm$ 2.40 21.23 $\pm$ 5.19
+B.D. 85.30 $\pm$ 6.24 69.87 $\pm$ 10.09 50.82 $\pm$ 11.74 10.57 $\pm$ 5.26 29.74 $\pm$ 11.28
+NICE 83.75 $\pm$ 3.21 65.82 $\pm$ 5.93 44.41 $\pm$ 6.12 7.04 $\pm$ 2.83 23.55 $\pm$ 4.86
+MLP 80.49 $\pm$ 2.61 62.09 $\pm$ 3.88 40.37 $\pm$ 3.71 5.67 $\pm$ 1.09 19.97 $\pm$ 3.02
+ATM-S (Ours) 87.36 $\pm$ 3.97 72.80 $\pm$ 7.02 53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
+ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51 53.15 $\pm$ 6.52 11.12 $\pm$ 3.26 32.61 $\pm$ 6.78
 Table 7: Accuracy of zero-shot Retrieval on THINGS-EEG for odd-numbered subjects(batch size=16).

### v5

@@
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet 83.23 $\pm$ 2.77 65.67 $\pm$ 3.84 43.87 $\pm$ 3.78 7.79 $\pm$ 1.27 22.12 $\pm$ 2.88
-EEGConformer 89.18 $\pm$ 2.88 75.49 $\pm$ 5.30 55.32 $\pm$ 6.88 11.29 $\pm$ 3.55 33.49 $\pm$ 6.78
-ShallowFBCSPNet 84.29 $\pm$ 2.56 67.13 $\pm$ 3.61 46.07 $\pm$ 4.55 8.26 $\pm$ 2.95 25.87 $\pm$ 5.03
-EEGNetV4 89.03 $\pm$ 2.86 75.40 $\pm$ 5.18 56.77 $\pm$ 6.60 13.29 $\pm$ 3.99 35.50 $\pm$ 7.14
-B.D. 91.14 $\pm$ 2.45 79.59 $\pm$ 4.59 62.62 $\pm$ 6.33 16.29 $\pm$ 4.35 42.16 $\pm$ 7.56
-NICE 92.17 $\pm$ 2.78 82.10 $\pm$ 5.55 65.65 $\pm$ 8.03 19.40 $\pm$ 5.83 46.26 $\pm$ 10.42
-MLP 87.67 $\pm$ 4.41 72.58 $\pm$ 7.62 53.52 $\pm$ 9.45 11.81 $\pm$ 4.33 32.74 $\pm$ 9.24
-ATM-S (Ours) 93.89 $\pm$ 1.97 85.38 $\pm$ 3.95 71.63 $\pm$ 6.17 22.84 $\pm$ 5.87 52.22 $\pm$ 8.34
+EEGITNet
+83.23 $\pm$ 2.77
+65.67 $\pm$ 3.84
+43.87 $\pm$ 3.78
+7.79 $\pm$ 1.27
+22.12 $\pm$ 2.88
+EEGConformer
+89.18 $\pm$ 2.88
+75.49 $\pm$ 5.30
+55.32 $\pm$ 6.88
+11.29 $\pm$ 3.55
+33.49 $\pm$ 6.78
+ShallowFBCSPNet
+84.29 $\pm$ 2.56
+67.13 $\pm$ 3.61
+46.07 $\pm$ 4.55
+8.26 $\pm$ 2.95
+25.87 $\pm$ 5.03
+EEGNetV4
+89.03 $\pm$ 2.86
+75.40 $\pm$ 5.18
+56.77 $\pm$ 6.60
+13.29 $\pm$ 3.99
+35.50 $\pm$ 7.14
+B.D.
+91.14 $\pm$ 2.45
+79.59 $\pm$ 4.59
+62.62 $\pm$ 6.33
+16.29 $\pm$ 4.35
+42.16 $\pm$ 7.56
+NICE
+92.17 $\pm$ 2.78
+82.10 $\pm$ 5.55
+65.65 $\pm$ 8.03
+19.40 $\pm$ 5.83
+46.26 $\pm$ 10.42
+MLP
+87.67 $\pm$ 4.41
+72.58 $\pm$ 7.62
+53.52 $\pm$ 9.45
+11.81 $\pm$ 4.33
+32.74 $\pm$ 9.24
+ATM-S (Ours)
+93.89 $\pm$ 1.97
+85.38 $\pm$ 3.95
+71.63 $\pm$ 6.17
+22.84 $\pm$ 5.87
+52.22 $\pm$ 8.34
 ATM-E (Ours) 94.44 $\pm$ 1.76 86.11 $\pm$ 3.39 71.89 $\pm$ 5.64 24.38 $\pm$ 6.56 54.14 $\pm$ 8.08
@@
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet 77.14 $\pm$ 4.14 55.29 $\pm$ 6.22 31.88 $\pm$ 6.20 2.93 $\pm$ 1.70 13.00 $\pm$ 4.45
-EEGConformer 79.49 $\pm$ 3.91 59.31 $\pm$ 5.83 36.67 $\pm$ 5.35 4.11 $\pm$ 1.83 17.04 $\pm$ 4.21
-ShallowFBCSPNet 75.14 $\pm$ 4.45 53.06 $\pm$ 6.70 31.23 $\pm$ 6.56 3.19 $\pm$ 2.17 13.01 $\pm$ 5.15
-EEGNetV4 82.60 $\pm$ 3.17 64.28 $\pm$ 5.44 42.24 $\pm$ 6.10 6.13 $\pm$ 2.40 21.23 $\pm$ 5.19
-B.D. 81.49 $\pm$ 3.52 62.35 $\pm$ 6.44 40.53 $\pm$ 7.19 6.16 $\pm$ 2.40 20.45 $\pm$ 5.40
-NICE 81.85 $\pm$ 2.53 63.57 $\pm$ 4.46 41.86 $\pm$ 4.68 6.43 $\pm$ 1.46 21.39 $\pm$ 3.62
-MLP 80.49 $\pm$ 2.61 62.09 $\pm$ 3.88 40.37 $\pm$ 3.71 5.67 $\pm$ 1.09 19.97 $\pm$ 3.02
-ATM-S (Ours) 82.88 $\pm$ 5.18 65.03 $\pm$ 8.51 44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
-ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98 44.84 $\pm$ 7.76 7.47 $\pm$ 2.84 23.75 $\pm$ 6.86
+EEGITNet
+77.14 $\pm$ 4.14
+55.29 $\pm$ 6.22
+31.88 $\pm$ 6.20
+2.93 $\pm$ 1.70
+13.00 $\pm$ 4.45
+EEGConformer
+79.49 $\pm$ 3.91
+59.31 $\pm$ 5.83
+36.67 $\pm$ 5.35
+4.11 $\pm$ 1.83
+17.04 $\pm$ 4.21
+ShallowFBCSPNet
+75.14 $\pm$ 4.45
+53.06 $\pm$ 6.70
+31.23 $\pm$ 6.56
+3.19 $\pm$ 2.17
+13.01 $\pm$ 5.15
+EEGNetV4
+82.60 $\pm$ 3.17
+64.28 $\pm$ 5.44
+42.24 $\pm$ 6.10
+6.13 $\pm$ 2.40
+21.23 $\pm$ 5.19
+B.D.
+81.49 $\pm$ 3.52
+62.35 $\pm$ 6.44
+40.53 $\pm$ 7.19
+6.16 $\pm$ 2.40
+20.45 $\pm$ 5.40
+NICE
+81.85 $\pm$ 2.53
+63.57 $\pm$ 4.46
+41.86 $\pm$ 4.68
+6.43 $\pm$ 1.46
+21.39 $\pm$ 3.62
+MLP
+80.49 $\pm$ 2.61
+62.09 $\pm$ 3.88
+40.37 $\pm$ 3.71
+5.67 $\pm$ 1.09
+19.97 $\pm$ 3.02
+ATM-S (Ours)
+82.88 $\pm$ 5.18
+65.03 $\pm$ 8.51
+44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
+ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98
+44.84 $\pm$ 7.76
+7.47 $\pm$ 2.84
+23.75 $\pm$ 6.86
 Table 6: Overall accuracy of zero-shot Retrieval on THINGS-EEG dataset.
@@
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet 76.69 $\pm$ 12.97 56.98 $\pm$ 16.31 36.35 $\pm$ 15.11 5.75 $\pm$ 3.62 18.14 $\pm$ 9.40
-EEGConformer 76.17 $\pm$ 13.13 56.29 $\pm$ 16.70 34.72 $\pm$ 14.79 3.98 $\pm$ 2.80 17.10 $\pm$ 9.21
-ShallowFBCSPNet 74.32 $\pm$ 12.14 53.97 $\pm$ 15.81 33.48 $\pm$ 14.35 6.10 $\pm$ 4.61 16.53 $\pm$ 9.94
-EEGNetV4 92.81 $\pm$ 2.22 83.15 $\pm$ 4.20 67.81 $\pm$ 6.11 19.51 $\pm$ 5.19 48.99 $\pm$ 6.75
-B.D. 78.42 $\pm$ 8.81 58.24 $\pm$ 12.13 37.97 $\pm$ 11.38 5.88 $\pm$ 3.49 18.61 $\pm$ 7.81
-NICE 92.73 $\pm$ 2.75 83.26 $\pm$ 5.47 67.96 $\pm$ 8.31 19.32 $\pm$ 5.33 49.26 $\pm$ 9.69
-MLP 83.09 $\pm$ 2.54 66.70 $\pm$ 4.16 45.43 $\pm$ 4.58 7.23 $\pm$ 1.66 25.14 $\pm$ 3.66
+EEGITNet
+76.69 $\pm$ 12.97
+56.98 $\pm$ 16.31
+36.35 $\pm$ 15.11
+5.75 $\pm$ 3.62
+18.14 $\pm$ 9.40
+EEGConformer
+76.17 $\pm$ 13.13
+56.29 $\pm$ 16.70
+34.72 $\pm$ 14.79
+3.98 $\pm$ 2.80
+17.10 $\pm$ 9.21
+ShallowFBCSPNet
+74.32 $\pm$ 12.14
+53.97 $\pm$ 15.81
+33.48 $\pm$ 14.35
+6.10 $\pm$ 4.61
+16.53 $\pm$ 9.94
+EEGNetV4
+92.81 $\pm$ 2.22
+83.15 $\pm$ 4.20
+67.81 $\pm$ 6.11
+19.51 $\pm$ 5.19
+48.99 $\pm$ 6.75
+B.D.
+78.42 $\pm$ 8.81
+58.24 $\pm$ 12.13
+37.97 $\pm$ 11.38
+5.88 $\pm$ 3.49
+18.61 $\pm$ 7.81
+NICE
+92.73 $\pm$ 2.75
+83.26 $\pm$ 5.47
+67.96 $\pm$ 8.31
+19.32 $\pm$ 5.33
+49.26 $\pm$ 9.69
+MLP
+83.09 $\pm$ 2.54
+66.70 $\pm$ 4.16
+45.43 $\pm$ 4.58
+7.23 $\pm$ 1.66
+25.14 $\pm$ 3.66
 ATM-S (Ours) 94.60 $\pm$ 1.93 86.88 $\pm$ 4.20 73.89 $\pm$ 5.93 26.09 $\pm$ 6.96 58.07 $\pm$ 8.16
-ATM-E (Ours) 92.99 $\pm$ 2.20 83.81 $\pm$ 4.46 68.87 $\pm$ 7.27 22.40 $\pm$ 6.62 50.59 $\pm$ 9.59
+ATM-E (Ours)
+92.99 $\pm$ 2.20
+83.81 $\pm$ 4.46
+68.87 $\pm$ 7.27
+22.40 $\pm$ 6.62
+50.59 $\pm$ 9.59
 Subject independent - leave one subject for test (batch size=1024)
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet 77.47 $\pm$ 3.75 55.75 $\pm$ 6.01 33.60 $\pm$ 6.16 3.80 $\pm$ 2.01 14.02 $\pm$ 3.89
-EEGConformer 68.02 $\pm$ 7.75 44.24 $\pm$ 9.25 23.84 $\pm$ 7.65 1.66 $\pm$ 1.37 8.62 $\pm$ 4.20
-ShallowFBCSPNet 76.40 $\pm$ 4.64 54.81 $\pm$ 7.09 32.37 $\pm$ 7.28 2.49 $\pm$ 1.57 13.24 $\pm$ 5.98
-EEGNetV4 82.60 $\pm$ 3.17 64.28 $\pm$ 5.44 42.24 $\pm$ 6.10 6.13 $\pm$ 2.40 21.23 $\pm$ 5.19
-B.D. 85.30 $\pm$ 6.24 69.87 $\pm$ 10.09 50.82 $\pm$ 11.74 10.57 $\pm$ 5.26 29.74 $\pm$ 11.28
-NICE 83.75 $\pm$ 3.21 65.82 $\pm$ 5.93 44.41 $\pm$ 6.12 7.04 $\pm$ 2.83 23.55 $\pm$ 4.86
-MLP 80.49 $\pm$ 2.61 62.09 $\pm$ 3.88 40.37 $\pm$ 3.71 5.67 $\pm$ 1.09 19.97 $\pm$ 3.02
-ATM-S (Ours) 87.36 $\pm$ 3.97 72.80 $\pm$ 7.02 53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
-ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51 53.15 $\pm$ 6.52 11.12 $\pm$ 3.26 32.61 $\pm$ 6.78
+EEGITNet
+77.47 $\pm$ 3.75
+55.75 $\pm$ 6.01
+33.60 $\pm$ 6.16
+3.80 $\pm$ 2.01
+14.02 $\pm$ 3.89
+EEGConformer
+68.02 $\pm$ 7.75
+44.24 $\pm$ 9.25
+23.84 $\pm$ 7.65
+1.66 $\pm$ 1.37
+8.62 $\pm$ 4.20
+ShallowFBCSPNet
+76.40 $\pm$ 4.64
+54.81 $\pm$ 7.09
+32.37 $\pm$ 7.28
+2.49 $\pm$ 1.57
+13.24 $\pm$ 5.98
+EEGNetV4
+82.60 $\pm$ 3.17
+64.28 $\pm$ 5.44
+42.24 $\pm$ 6.10
+6.13 $\pm$ 2.40
+21.23 $\pm$ 5.19
+B.D.
+85.30 $\pm$ 6.24
+69.87 $\pm$ 10.09
+50.82 $\pm$ 11.74
+10.57 $\pm$ 5.26
+29.74 $\pm$ 11.28
+NICE
+83.75 $\pm$ 3.21
+65.82 $\pm$ 5.93
+44.41 $\pm$ 6.12
+7.04 $\pm$ 2.83
+23.55 $\pm$ 4.86
+MLP
+80.49 $\pm$ 2.61
+62.09 $\pm$ 3.88
+40.37 $\pm$ 3.71
+5.67 $\pm$ 1.09
+19.97 $\pm$ 3.02
+ATM-S (Ours)
+87.36 $\pm$ 3.97
+72.80 $\pm$ 7.02
+53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
+ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51
+53.15 $\pm$ 6.52
+11.12 $\pm$ 3.26
+32.61 $\pm$ 6.78
 Table 7: Accuracy of zero-shot Retrieval on THINGS-EEG for odd-numbered subjects(batch size=16).

### v6

@@
-Shenzhen, China
+Visual Decoding and Reconstruction via EEG Embeddings with Guided Diffusion
 Abstract
 How to decode human vision through neural signals has attracted a long-standing interest in neuroscience and machine learning.
-Modern contrastive learning and generative models improved the performance of fMRI-based visual decoding and reconstruction.
-However, the high cost and low temporal resolution of fMRI limit their applications in brain-computer interfaces (BCIs), prompting a high need for EEG-based visual reconstruction.
-In this study, we present an EEG-based visual reconstruction framework.
-It consists of a plug-and-play EEG encoder called the Adaptive Thinking Mapper (ATM), which is aligned with image embeddings, and a two-stage EEG guidance image generator that first transforms EEG features into image priors and then reconstructs the visual stimuli with a pre-trained image generator.
-Our approach allows EEG embeddings to achieve superior performance in image classification and retrieval tasks.
-Our two-stage image generation strategy vividly reconstructs images seen by humans.
-Furthermore, we analyzed the impact of signals from different time windows and brain regions on decoding and reconstruction.
+Modern contrastive learning and generative models improved the performance of visual decoding and reconstruction based on functional Magnetic Resonance Imaging (fMRI).
+However, the high cost and low temporal resolution of fMRI limit their applications in brain-computer interfaces (BCIs), prompting a high need for visual decoding based on electroencephalography (EEG).
+In this study, we present an end-to-end EEG-based visual reconstruction zero-shot framework, consisting of a tailored brain encoder, called the Adaptive Thinking Mapper (ATM), which projects neural signals from different sources into the shared subspace as the clip embedding, and a two-stage multi-pipe EEG-to-image generation strategy.
+In stage one, EEG is embedded to align the high-level clip embedding, and then the prior diffusion model refines EEG embedding into image priors.
+A blurry image also decoded from EEG for maintaining the low-level feature.
+In stage two, we input both the high-level clip embedding, the blurry image and caption from EEG latent to a pre-trained diffusion model.
+Furthermore, we analyzed the impacts of different time windows and brain regions on decoding and reconstruction.
 The versatility of our framework is demonstrated in the magnetoencephalogram (MEG) data modality.
-We report that EEG-based visual decoding achieves SOTA performance, highlighting the portability, low cost, and high temporal resolution of EEG, enabling a wide range of BCI applications.
-Our code is available at https://github.com/dongyangli-del/EEG_Image_decode.
-Keywords:
-EEG Visual reconstruction Neural alignment MEG
-[…]
-Figure 1: EEG/MEG-based visual retrieval, classification, and reconstruction tasks.
-(a) Overview of three visual decoding tasks using EEG/MEG data under natural image stimuli.
-(b) Comparisons of nine encoders on the THINGS-EEG dataset, including within-subject and cross-subject performance.
-(c) Comparisons on the THINGS-MEG dataset, similar to (b).
-Our ATM-S encoder achieves the highest performance compared to other competing encoders in EEG/MEG-based visual decoding tasks.
+The experimental results indicate that our EEG-based visual zero-shot framework achieves SOTA performance in classification, retrieval and reconstruction, highlighting the portability, low cost, and high temporal resolution of EEG, enabling a wide range of BCI applications.
+Our code is available at https://github.com/ncclab-sustech/EEG_Image_decode.
+Figure 1: EEG/MEG-based zero-shot brain decoding and reconstruction. Left: Overview of three visual decoding tasks using EEG/MEG data under natural image stimulus. Right: Our reconstruction examples.
 1 Introduction
-A key technical challenge in brain-computer interfaces (BCIs) is to decode/reconstruct the visual world seen by humans through non-invasive brain recordings, such as functional magnetic resonance imaging (fMRI), magnetoencephalography (MEG) or electroencephalography (EEG).
+A key technical challenge in BCIs is to decode/reconstruct the visual world seen by humans through non-invasive brain recordings, such as fMRI, MEG or EEG.
 These highly dynamic brain activities reflect human perception of the visual world, which is influenced by properties of the external visual stimulus, our internal states, emotions and even personal experiences.
-Thus, visual decoding and reconstruction based on neural signals can uncover how the human brain processes and interprets natural visual stimuli, as well as promote non-invasive BCI applications.
-Figure 2: EEG/MEG-based visual decoding and generation framework.
-The EEG encoder is designed as a flexible replacement component.
-After aligning with image features, the EEG features are used for zero-shot retrieval and classification tasks, and the reconstructed images are obtained through a two-stage generator.
+Thus, visual decoding and reconstruction based on neural signals can uncover how the human brain processes and interprets natural visual stimulus, as well as promote non-invasive BCI applications.
 Contrastive learning and generative models have greatly advanced fMRI-based visual decoding in both decoding tasks (e.g., image classification and retrieval) and generative tasks (e.g., image reconstruction).
-By combining pre-trained visual models, existing fMRI decoding models can learn highly-refined feature embeddings in limited data [28, 40].
-Using these embedded fMRI features, generative models such as diffusion models can reconstruct the image one is seeing [40, 12].
+By combining pre-trained visual models, existing fMRI decoding models can learn highly-refined feature embeddings in limited data Liu et al., 2023b ; Scotti et al., (2023).
+Using these embedded fMRI features, generative models such as diffusion models can reconstruct the image one is seeing Scotti et al., (2023); Fang et al., (2023).
 However, despite many advances in fMRI-based visual decoding, fMRI equipment is unportable, expensive, and difficult to operate, largely limiting its application in BCIs.
 Alternatively, EEG is portable, cheap, and universal, facilitating a wide range of BCI applications.
-EEG has higher temporal resolution and can effectively capture rapid changes in brain activity when processing complex, dynamic visual stimuli.
-However, EEG has long been considered incomparable to fMRI in natural image decoding/reconstruction tasks, as EEG suffers from low signal-to-noise ratio, low spatial resolution, and large inter-subject variability.
-Recent advances in multimodal alignment have made MEG/EEG visual decoding possible, although the performance is still inferior to fMRI [8, 46, 16].
+EEG has higher temporal resolution and can effectively capture rapid changes in brain activity when processing complex, dynamic visual stimulus.
+EEG has long been considered incomparable to fMRI in natural image decoding/reconstruction tasks, as EEG suffers from low signal-to-noise ratio, low spatial resolution, and large inter-subject variability.
+Recent advances in multimodal alignment have made MEG/EEG visual decoding possible, although the performance is still inferior to fMRI Cichy and Pantazis, (2017); Song et al., (2023); Grootswagers et al., (2022).
 Yohann Benchetrit et al.
@@
 Recently, Song et al.
-[46] used an EEG encoder based on ShallowNet [37] and performed representation alignment through contrastive learning, achieving excellent decoding performance on the THING-EEG dataset [14].
+Song et al., (2023) used an EEG encoder based on ShallowNet Schirrmeister et al., (2017) and performed representation alignment through contrastive learning, achieving excellent decoding performance on the THING-EEG dataset Gifford et al., (2022).
 These two studies provide preliminary evidence of the potential of EEG/MEG-based visual decoding.
 However, there is a significant gap in their performance compared to the fMRI-level performance.
-This gap is largely caused by weaknesses in its EEG encoder, which prevents EEG embeddings from effectively aligning with image embeddings.
-To fill in the gap, we develop an EEG/MEG-based visual decoding framework, including a novel EEG encoder, called Adaptive Thinking Mapper (ATM), and a two-stage image generation strategy.
-We survey existing EEG encoder modules, such as EEGNetV4  [25], ShallowNet [37], and Conformer  [44], and demonstrate through extensive comparative experiments and ablation studies that our method achieves state-of-the-art (SOTA) performance on image classification, retrieval, and generation tasks.
-Our work has three main contributions:
-  • We employ the attention module and a spatiotemporal convolution module in EEG encoder, which can be used in a plug-and-play manner, to extract representations in real-time for both EEG data and MEG data.
-  • Our brain encoder is trained with self-supervised contrastive learning framework, achieving SOTA performance on EEG and MEG datasets from natural image decoding tasks.
-  • We present a two-stage image generation strategy to facilitate the adjustment of the distribution from the original EEG data to the corresponding images.
-It employs a prior diffusion model that is conditioned on the EEG embeddings to generate the image embeddings, followed an enhanced Stable Diffusion module to generate images with EEG priors.
-Figure 3: The structure of EEG encoder ATM.
-According to the position of EEG channels and the spatiotemporal property of EEG signals, we design the position encoding and temporospatial encoding in ATM.
+This gap is largely caused by the fact that the framework of EEG visual decoding and reconstruction have not yet been thoroughly explored.
+To fill this gap, we have developed a visual decoding framework based on EEG/MEG, including a novel EEG encoder and a two-stage image generation strategy. Our work has three main contributions:
+  1. 1.
+    We present brain decoding framework, which is the first work allows zero-shot image classification, retrieval, and reconstruction via EEG data.
+Experimental results demonstrate that our framework is applicable to various common EEG encoder architectures.
+  2. 2.
+    By extensively studying the existing EEG encoder modules, we construct a tailored EEG encoder ATM, which achieves state-of-the-art performance in three downstream visual decoding tasks.
+  3. 3.
+    We report a two-stage EEG-to-image generation strategy, which separately extracts high-level and low-level visual features from EEG and refining these features with an additional lightweight prior diffusion model, enabling reliable reconstruction of images using less than 500ms EEG.
+Figure 2: EEG/MEG-based visual decoding and generation framework.
+The EEG encoder is designed as a flexible replacement component.
+After aligning with image features, the EEG features are used for zero-shot retrieval and classification tasks, and the reconstructed images are obtained through a two-stage generator.
 2 Method
-To learn high-quality latent representations of EEG data, it is crucial to consider the spatial position of EEG channels and the spatiotemporal properties of EEG signals.
+To learn high-quality latent representations of EEG data, it is crucial to consider the spatial position of EEG channels and the Temporal-Spatial properties of EEG signals.
 Let $T$ represent the length of the time window of the data, $C$ the number of EEG channels, and $N$ the total number of data samples.
-Our objective is to derive EEG embeddings $Z_{E}=f(E)\in\mathbb{R}^{N\times F}$ from the brain activity data $E\in\mathbb{R}^{N\times C\times T}$, where $f$ is the EEG encoder and $F$ is the feature dimension of the embeddings.
+Our objective is to derive EEG embeddings $Z_{E}=f(E)\in\mathbb{R}^{N\times F}$ from the brain activity data $E\in\mathbb{R}^{N\times C\times T}$, where $f$ is the EEG encoder and $F$ is the projection dimension of the embeddings.
 Concurrently, we use the CLIP model to extract image embeddings $Z_{I}\in\mathbb{R}^{N\times F}$ from images $I$.
@@
 In the training phase, the EEG encoder is trained with EEG and image pairs using a contrastive learning framework.
-In the inference phase, the EEG embeddings from the trained plug-and-play EEG encoder can be used for a variety of tasks, including EEG-based image classification, retrieval, and EEG-guided image generation.
+In the inference phase, the EEG embeddings from the trained EEG projector can be used for a variety of zero-shot tasks, including EEG-based image classification, retrieval, and image reconstruction.
 2.1 ATM for EEG Embedding
-We develop an EEG encoder, called Adaptive Thinking Mapper (ATM), for aligning the original EEG signals to its feature representation space (Fig.
+Inspired by advanced time series models Liu et al., 2023a ; Gao et al., (2024), we develop an EEG encoder called ATM, for aligning the original EEG signals to its feature representation space (Fig.
 3).
-ATM is based on the Transformer Encoder and spatiotemporal convolution architecture.
-Specifically, we utilize a one-dimensional linear layer to project the input one-dimensional EEG data to the embedding dimension required by the Transformer model.
-These inputs are processed through a self-attention module to integrate the embeddings of input data and positional encoding.
-We first divide the input into fixed-size patches, with each image patch containing a fixed number of data points.
-Then, each patch is mapped to a high-dimensional embedding space through a linear layer.
-Subsequently, through the temporal aggregation process, we obtain the processed embedding representations.
-Notably, ATM addresses the inadequacies of other modules in modeling capabilities on the temporal scale through the Input Layer.
-For it has been reported that the spatiotemporal convolution module with a large convolution kernel is an effective way to represent EEG data with a small number of parameters [46], we use a similar spatiotemporal convolution (STConv)-based feature extraction module.
-The difference is our STConv Module is plug-and-play and can be flexibly replaced with different types of spatiotemporal convolution components as needed to adapt to various EEG/MEG datasets.
-Finally, MLP projection layer consists of $M$ simple residual components and fully connected layers, with LayerNorm applied in the output to ensure the stability of training.
+ATM is based on the channel-wise Transformer encoder, Temporal-Spatial convolution and multilayer perceptron (MLP) architecture.
+In contrast to other conventional practices, the original EEG does not need to be segmented, and each sequence acts as a patch.
+After sinusoidal position embedding, these patches are processed through a channel-attention module to integrate the information of different series.
+Subsequently, through the Temporal-Spatial aggregation, we project the output with a MLP to get rational shape representations.
+The Temporal-Spatial convolution module is an effective way to represent EEG data with a small number of parameters (Song et al., , 2023), prevent overfitting in training.
+The difference is our components is plug-and-play and can be flexibly replaced with different types of Temporal-Spatial convolution components as needed to adapt to various EEG/MEG datasets.
+Finally, MLP projectior consists of $M$ simple residual components and fully connected layers, with LayerNorm applied in the output to ensure the stability of training.
+In addition to entering the original series, we provide an identification input for a known subject and can specifically use this token for downstream tasks.
+For unknown subjects, we use shared tokens or average all tokens equally directly into the MLP projector.
+Figure 3: The structure of ATM.
+The original EEG sequences of different variates are independently embedded into tokens.
+Channel-wise attention is applied to embedded variate tokens with enhanced interpretability revealing electrode correlations.
+And representations of each token are extracted by the shared feedforward network (FFN).
+Then Temporal-Spatial convolution can prevent overfitting and enhance the ability of Temporal-Spatial modeling.
 2.2 Image Embedding
-Many previous studies have explored various training strategies to train deep neural networks for image embedding, such as VGG-19 and ResNet trained with supervised learning, CLIP trained with contrastive learning, and VAEs with self-supervised learning [49, 4, 46, 16].
-They have reported that CLIP models pre-trained using the Vision Transformer (ViT) architecture perform better in a range of downstream tasks, including image decoding and reconstruction, compared to models trained using supervised learning methods (such as VGG, ResNet) and self-supervised VAE frameworks.
-Thus, in this study, we use CLIP for image embedding and align the CLIP’s output image representation, denoted as $Z_{I}\in\mathbb{R}^{N\times 1024}$, with the EEG embeddings.
-Before formal training, all images undergo standard preprocessing  [34].
+Many previous studies have explored various training strategies to train deep neural networks for image embedding, such as VGG-19 and ResNet trained with supervised learning, CLIP, DINO trained with contrastive learning, and VAEs with self-supervised learning Teichmann et al., (2023); Benchetrit Y, (2023); Song et al., (2023); Grootswagers et al., (2022).
+They have reported that DINO and CLIP models pre-trained using the Vision Transformer (ViT) architecture perform better in a range of downstream tasks, including image decoding and reconstruction, compared to models trained using supervised learning methods (such as VGG, ResNet) and self-supervised VAE frameworks.
+Thus, in this study, we use CLIP for image embedding, denoted as $Z_{I}\in\mathbb{R}^{N\times 1024}$, instead of $Z_{I}\in\mathbb{R}^{N\times 257\times 768}$, with the EEG embeddings.
+Before formal training, all images undergo the standard preprocessing procedure Radford et al., (2021).
 2.3 EEG Guidance Image Generation
-In this study, we present a two-stage pipeline for generating images that serve as visual stimuli for EEG recordings, as shown in the bottom right of Fig.
+In this study, we present a two-stage pipeline for generating images that serve as visual stimulus for EEG recordings, as shown in the bottom right of Fig.
 2.
@@
 In Stage I, we first focus on the prior diffusion stage.
-Inspired by DALL-E 2 [35] and Mind’s Eyes [40], we train a diffusion model conditioned on the EEG embeddings $\hat{Z}_{E}$ to learn the distribution of CLIP embeddings $p(z_{I}|z_{E})$.
+Inspired by DALL-E 2 Ramesh et al., (2022) and Mind’s Eyes Scotti et al., (2023), we train a diffusion model conditioned on the EEG embeddings $\hat{Z}_{E}$ to learn the distribution of CLIP embeddings $p(z_{I}|z_{E})$.
 In this stage, we construct a lightweight U-Net: $\epsilon_{\text{prior}}(z^{t}_{I},t,z_{E})$, where $z^{t}_{I}$ represents the noisy CLIP embedding at diffusion time step $t$.
@@
 Through this diffusion model, we can generate corresponding CLIP embeddings $z_{I}$ from EEG embeddings as a prior for stage II.
-In Stage II, we employ the pre-trained SDXL [33] and IP-Adapter [59] models to model the generator $p(I|z_{I})$, thereby sampling image $I$ according to $z_{I}$.
-This stage mainly focuses on converting CLIP embeddings into corresponding images.
-Further details are provided in Appendix 0.C.
+In Stage II, we employ the pre-trained SDXL Podell et al., (2023) and IP-Adapter Ye et al., (2023) models to model the generator $p(I|z_{I})$, thereby sampling image $I$ according to $z_{I}$.
+In addition, we introduce the low-level features here using img2imgMeng et al., (2021).
+Further details are provided in Appendix C.
 2.4 Loss Function
 Following the methodology outlined by Benchetrit et al.
-[4], we adopt a dual approach to loss functions, serving distinct objectives.
+Benchetrit Y, (2023), we adopt a dual approach to loss functions, serving distinct objectives.
 For the classification and retrieval tasks, we only utilize the CLIP loss, which is inspired by the contrastive learning approach described in Radford et al.
-[34].
-This loss function aids in aligning the EEG data $E$ with corresponding image data $I$, thereby facilitating the identification of EEG-image pairs and maximizing the EEG classification boundaries.
-For the generation tasks, besides the CLIP loss, we add a Mean Squared Error (MSE) loss to facilitate learning in regression scenarios.
+Radford et al., (2021).
+This loss function aids in aligning the EEG data $E$ with corresponding image data $I$, thereby facilitating the identification of EEG-image pairs and maximizing the boundaries of EEG representations.
+For the generation tasks, besides the CLIP loss, we add a Mean Squared Error (MSE) loss to facilitate consistency learning in regression.
 Thus the overall loss function for our model is a combination of these two distinct loss types, expressed as:
@@
 Here, $\lambda$ is a hyperparameter that balances the contribution of each loss type.
-Figure 4: EEG-based image classification. (a) The paradigm of EEG-based image classification. (b) Average accuracy across different methods in the subjects.
-Figure 5: EEG-based image retrieval.
-(a) The paradigm of EEG-based image retrieval.
-(b) Average accuracy across different methods in the subjects.
-(c) Images showing the top-5 accuracy in EEG-image retrieval tasks.
-See Appendix 0.F for additional images results.
-(d) Representation similarity analysis.
-Clustering of image features and EEG features using a k-means algorithm with k=5 is performed, followed by the calculation of similarity between clusters.
-See Appendix 0.E for additional Representational similarity results.
 3 Experiments
 3.1 Training and Computational Considerations
-We conducted our experiments on the THINGS-EEG dataset’s training set [14, 16].
-To verify the versatility of ATM for embedding electrophysiological data, we tested it on MEG data modality using the THINGS-MEG dataset [18].
-We used the Adam optimizer [24] to train the across-subject model on a set of approximately 496,200 samples, and the within-subject model on a set of about 66,160 samples, with a learning rate of $3\times 10^{-4}$ and batch sizes of 16 and 1024.
+We conducted our experiments on the THINGS-EEG dataset’s training set Gifford et al., (2022); Grootswagers et al., (2022).
+To verify the versatility of ATM for embedding electrophysiological data, we tested it on MEG data modality using the THINGS-MEG dataset Hebart et al., (2023).
+All experiments can be completed in a single NVIDIA RTX 4090 GPU.
+We used the Adam optimizer Kingma and Ba, (2014) to train the across-subject model on a set of approximately 496,200 samples, and the within-subject model on a set of about 66,160 samples, with an initial learning rate of $3\times 10^{-4}$ and batch sizes of 16 and 1024.
 Our initial temperature parameter was set to 0.07.
@@
 For fairness, all models’ hyperparameters were kept consistent.
-In our study, we compared the performance of different encoders on the within-subject test set and cross-subject (leave-one-subject-out) test set (see Appendix 0.G).
+In our study, we compared the performance of different encoders on the within-subject test set and cross-subject (leave-one-subject-out) test set (see Appendix H).
+Figure 4: EEG/MEG-based decoding and reconstruction performance.
+Left: Comparisons of nine encoders on the THINGS-EEG dataset, including within-subject and cross-subject performance.
+Right: Comparisons on the THINGS-MEG dataset, similar to left.
+Our method achieves the highest performance compared to other competing encoders in EEG/MEG-based visual decoding tasks.
 3.2 EEG Decoding Performance
-The plug-and-play ATM can obtain the EEG embedding for the classification task using a simple cosine similarity measurement.
-We output the category with the highest cosine similarity (Fig.
-4a).
-More details of the EEG-based image classification are in Appendix 0.B.
+Our method obtains the EEG embedding for the classification task.
+We output the category of EEG with the highest cosine similarity with text embeddings(Fig.
+5a).
 Fig.
-4b presents the average accuracy across different methods in the subjects, and shows our method outperforms others.
-Here, we test the effectiveness of ATM-extracted EEG embeddings in the image retrieval task.
+5c presents the average accuracy across different methods in the subjects, and shows that our method outperforms others.
+More details of the EEG-based image classification are in Appendix B.
+In Fig.
+5, we test the effectiveness of EEG embeddings in the image retrieval task.
+We calculate the cosine similarity between the EEG embeddings and the CLIP embeddings instead of text embeddings in the image dataset (with 200 images).
 Fig.
-5a shows the image retrieval process.
-We calculate the cosine similarity between the extracted EEG embeddings and the CLIP embeddings of the image dataset (with 200 images), and output the image with the highest similarity as the retrieved image.
+5d shows the average results for all subjects.
+We take the highest test accuracy in the training process as the statistical result.
 Fig.
-5b shows the average results for all subjects in the in-subject test.
-We take the highest test accuracy in the training process as the statistical result.
-See the Appendix 0.G for more detailed averages of test accuracy in subjects.
+5b shows the Top-5 retrieved images corresponding to the real visual stimulus seen by subjects.
+Compared with the previous models, the Top-1 accuracy of our model is significantly improved, and the Top-5 images all maintain a high degree of similarity with the original images.
+See the Tab.
+8 in Appendix for more detailed averages of test accuracy in subjects.
+Figure 5: EEG-based image retrieval and classification.
+(a) The paradigm of EEG-based image retrieval and classification.
+(b) Samples of the top-5 accuracy in EEG-image retrieval tasks.
+See Appendix G for additional images results.
+(c) Average in-subject classification accuracy across different methods.
+(d) Average in-subject retrieval accuracy across different methods.
+Ablation Study on ATM
+We systematically deconstructed and analyzed each layer of our EEG projector.
+We conducted an ablation study for each component in ATM (i.e., the MLP projector, the Temporal-Spatial convolution module and the channel-wise attention block).
+We specified two different convolution architectures, ShallowNet (ATM-S) and EEGNetV4 (ATM-E), as our convolution backbone.
+Appendix B.3 showed the results obtained under different experimental configurations.
+3.3 Image Generation Performance
 Fig.
-5c shows the Top-5 retrieved images corresponding to the real visual stimuli seen by subjects.
-Compared with the previous model, the Top-1 accuracy of our model is significantly improved, and the Top-5 images all maintain a high degree of similarity with the original images.
-We calculated the similarity of the EEG to the semantics and images.
-The similarity between the image and EEG is shown in Fig.
-5d.
-We performed a representational similarity analysis, and we could observe different degrees of intra-class aggregation.
-This means that the EEG representation obtained is closer to the representation of the corresponding image itself than to the semantics.
+6a shows the process of generating images under the guidance of EEG embedding and evaluating the quality of the generated images.
+To evaluate the generation performance, we conducted an image retrieval task.
+Specifically, we extract the CLIP embedding of the generated images and compare the similarity between the CLIP embeddings of all images to retrieve the generated image.
+Fig.
+6b shows the similarity of distribution.
+Fig.
+6c shows the generated samples.
+The generated images have high semantic similarity with the seen images and have good diversity in low-level visual features, which can be manipulated by the guidance scale hyperparameter (Fig.
+6d).
+We also report the decoding and reconstruction performance for EEG, MEG, and fMRI across various metrics from different datasets in the Tab.
+1.
 Figure 6: EEG guidance image generation.
-(a) The paradigm of EEG guidance image generation.
+(a) The paradigm of image generation.
 (b) The similarity between random visual objects and the EEG embeddings, and the similarity between generated visual objects and the target EEG embeddings.
 (c) Comparison between the original image and the image generated using the corresponding EEG data.
-(see Appendix 0.C for details).
+(see Appendix C for details).
 (d) The similarity between visual objects and target EEG embeddings as the guidance scale changes, and the diversity of visual objects as the guidance scale changes.
-See Appendix 0.F for additional results.
-Ablation study on ATM
-We systematically deconstructed and analyzed each layer of our EEG encoder module.
-We conducted an ablation study for each component in ATM (i.e., the MLP projector, the adaptive convolution module, the Input layer, and the spatial attention block).
-Notably, ATM is a flexible, plug-and-play module, which can be easily replaced with any form of convolutional component or even omitted.
-Appendix 0.B.3 showed that the adaptive convolutional components significantly enhanced encoder performance, while the Input layer reduced the model’s training consumption and substantially improved the existing components’ capabilities.
-Surprisingly, our spatial attention module based on sine-cosine positional encoding improved encoder performance only when used alone, but the benefits were not as significant when combined with the Input layer.
-3.3 Image Generation performance
-Using the image generation task, we verify the effectiveness of ATM’s EEG embedding and our two-stage generation model.
-Fig.
-6a shows the process of generating images under the guidance of EEG embedding and evaluating the quality of the generated images.
-For one test image, we put EEG embedding of the subject viewing the image into the two-stage generator to guide the generation of the image.
-To evaluate the generation performance, we conducted an image retrieval task.
-Specifically, we extract the CLIP embedding of the generated images and calculate the similarity between the CLIP embeddings of all images in the image dataset to retrieve the generated image.
-Figure 7: Examples of EEG-guided visual reconstruction.
-From top to bottom, we exhibit the best, median, and worst 12 generated images, respectively.
-We show the images subjects seen and the generated images by our two-stage image generator.
-See Appendix 0.F for additional results.
-Fig.
-6b shows the generation effect of the sample image.
-Fig.
-6c shows the retrieval accuracy using generated images.
-The generated images have high semantic similarity with the seen images and have good diversity in low-level visual features, which can be manipulated by the guidance scale hyperparameter (Fig.
-6d).
-Surprisingly, the retrieval accuracy using generated images is even higher than that using EEG embeddings, implying the benefit of our two-stage visual reconstruction for image retrieval (See Appendix 0.C for Fig.11).
-We also report the decoding and reconstruction performance for EEG, MEG, and fMRI across various metrics in the Appendix 0.D.
-We visualize the best, medium and worst generated images in Fig.
-7.
-We randomly selected the EEG data of a subject viewing 100 images, and extracted EEG embeddings to guide image generation.
-By calculating the cosine similarity of the CLIP embedding between the generated image and the original image, we found 12 images each with the best, medium and worst generation effects.
-It can be seen that in the best group, the generated image is not only highly consistent with the semantics of the original image, but also well retains the low-level visual features.
-in the medium group, the generated image maintains the semantic features of the original image, and the low-level visual features are well preserved.
-Visual features were altered.
-in the worst group, both semantic features and low-level visual features were altered.
+See Appendix G for additional results.
+Table 1: Quantitative assessments of the reconstruction quality for EEG, MEG, and fMRI in Subject 8. For detailed explanations of the metrics.
+Low-level High-level
+Dataset $\uparrow$ PixCorr $\uparrow$ SSIM $\uparrow$ AlexNet(2) $\uparrow$ AlexNet(5) $\uparrow$ Inception $\uparrow$ CLIP $\uparrow$ SwAV $\downarrow$
+NSD-fMRI Benchetrit Y, (2023) 0.305 0.366 0.962 0.977 0.910 0.917 0.410
+NSD-fMRI Ozcelik and VanRullen, (2023) 0.254 0.356 0.942 0.962 0.872 0.915 0.423
+NSD-fMRI Scotti et al., (2024) 0.130 0.308 0.917 0.974 0.936 0.942 0.369
+[…]
+THINGS-MEG (averaged) Benchetrit Y, (2023) 0.076 0.336 0.736 0.826 0.671 0.767 0.584
+THINGS-MEG (Ours) 0.104 0.340 0.613 0.672 0.619 0.603 0.651
+THINGS-EEG (Ours) 0.160 0.345 0.776 0.866 0.734 0.786 0.582
 3.4 Temporal Analysis
-To investigate the effects of EEG time window on visual decoding accuracy, we calculated the average top-1 classification accuracy for two different time windows: [0, $t$], including the entire period from the onset of visual stimuli to time point $t$, and [$t$-100, $t$], only including the data 100ms before time point $t$.
-We compared the accuracy with a randomly selected baseline (0.5% chance level) to test non-random predictive performance (Fig.
-8).
-Our results show that within 500ms after visual stimulus presentation, the EEG signal decoding accuracy reaches an upper limit of about 30%, after which the accuracy no longer improves (Fig.
-8a).
+To investigate the effects of different EEG time window on visual decoding, we calculated the average top-1 classification accuracy for sliding and growing time windows: [0, $t$], including the entire period from the onset of visual stimulus to time point $t$, and [$t$-100, $t$], only including the data 100ms before time point $t$.
+We compared the accuracy with a randomly selected baseline (0.5% chance level) to test predictive performance (Fig.
+7).
+Our results show that within 500ms after visual stimulus, the accuracy reaches an upper limit of about 30%, after which the accuracy no longer improves (Fig.
+7a).
 The MEG decoding shows a similar profile as the time window expands (Fig.
-8b).
+7b).
 We exhibit the generated images under different EEG time windows, [0, $t$] in Fig.
-8c.
-The similarity between the generated images and the original images is low when the time window is less than 150ms, then this similarity gradually increase as the time window expands.
-After 500 milliseconds, EEG-guided image generation can reliably reveal the semantics of the images seen, such as aircraft carriers, modems, jelly beans, et al.
+7c.
+The similarity is low when the time window is less than 150ms, and this similarity gradually increase as the time window expands.
+After 500 milliseconds, EEG-guided image generation can reliably reveal the semantics of the images seen.
 Interestingly, we find differences in the optimal reconstruction time windows for different categories of images, for example, jelly beans (200ms) are faster than aircraft carrier (500ms), implying that the human brain may process different visual objects at different speeds.
 This finding highlights the advantage of EEG’s high temporal resolution in studying fast visual processing compared with the lower temporal resolution of fMRI.
-Figure 8: Effects of different EEG/MEG time windows on EEG-guided visual retrieval and reconstruction.
+Figure 7: Performance of different EEG/MEG time windows on EEG-guided visual retrieval and reconstruction.
 (a) The retrieval accuracy of the expanding EEG windows at intervals [0, t] and at intervals [t-100, t] respectively.
 (b) The retrieval accuracy of the expanding MEG windows.
-(c) Images reconstructed by EEG as the EEG window expands.
-When the EEG time window is greater than 200ms, the reconstructed image is reliable.
+(c) Samples reconstructed as the EEG window expands.
+When the EEG time window is greater than 200ms, the reconstructed image is stable.
+See Appendix for more detailed explanation.
 3.5 Spatial Analysis
-To examine the contribution of different brain regions to visual decoding, we divided the EEG electrodes from the THING-EEG data into five distinct brain regions (i.e., Frontal, Temporal, Center, Parietal, Occipital regions in Fig.
-9a), and then conducted ablation experiments on retrieval task (Fig.
-9b) and the reconstruction task (Fig.
-9c).
+To investigate the contribution of different brain regions to visual decoding, we divided the EEG electrodes from the THING-EEG data into five distinct brain regions (i.e., Frontal, Temporal, Center, Parietal, Occipital regions in Fig.
+8a), and then conducted ablation experiments on retrieval task (Fig.
+8b) and the reconstruction task (Fig.
+8c).
 The results showed that using information from all brain regions is optimal, for both retrieval and generation tasks.
@@
 Parietal and temporal regions contain some semantic information, whereas frontal and central regions contribute the least useful information to the visual decoding.
-Figure 9: EEG-guided visual retrieval and reconstruction using EEG signals from different brain regions.
+Figure 8: EEG-guided retrieval and reconstruction using EEG from different brain regions.
 (a) The EEG electrodes assigned to five brain regions.
@@
 4 Related Works
-Visual decoding using neural signal: Decoding visual information from our brain has been a long-standing pursuit in neuroscience and computer science [30, 20].
-Some progress has been made in decoding steady-state visual stimuli.
-However, accurately and rapidly decoding semantic information in natural images remains a challenge [42].
-fMRI has been widely used to estimate semantic and shape information in visual processing within the brain [48, 19].
+Visual decoding using neural signals: Decoding visual information from our brain has been a long-standing pursuit in neuroscience and computer science Miyawaki et al., (2008); Jia et al., (2021).
+Some progress has been made in decoding steady-state visual stimulus.
+However, accurately and rapidly decoding semantic information in natural images remains a challenge Shi et al., (2023).
+fMRI has been widely used to estimate semantic and shape information in visual processing within the brain Takagi and Nishimoto, (2023); Ho et al., (2023).
 However, the demand for high-speed and practical applications in brain-computer interfaces calls for alternative approaches.
-EEG, due to its high temporal resolution and portability, emerges as a promising option [53].
-Yet, the overall performance across different subjects and biological plausibility remains unresolved [1].
-Furthermore, previous approaches often relied on supervised learning methods with limited image categories, overlooking the intrinsic relationship between image stimuli and brain responses [28, 41, 27].
-Neural decoding for EEG signal: Previous studies have shown the efficacy of spatiotemporal modules in representing neural data [37, 25].
-For example, lightweight convolutional neural networks such as EEGNet and ShallowNet  [37] have achieved considerable performance in small EEG and MEG datasets.
-Using contrastive learning, it has been shown that merely using convolutional neural networks and projection layers can yield satisfactory results on neural datasets [6].
-More recently, Benchetrit et al proposed a B.D.
-encoder for MEG embedding, allowing real-time MEG-based reconstruction of visual perception [4].
+EEG, due to its high temporal resolution and portability, emerges as a promising option Willett et al., (2021).
+Yet, the overall performance across different subjects and biological plausibility remains unresolved Ahmed et al., (2021).
+Furthermore, previous approaches often relied on supervised learning methods with limited image categories, overlooking the intrinsic relationship between image stimulus and brain responses Liu et al., 2023b ; Shen et al., (2019); Li et al., (2020).
+Neural decoding using EEG/MEG data: Previous studies have shown the efficacy of Temporal-Spatial modules in representing neural data Schirrmeister et al., (2017); Lawhern et al., (2018).
+For example, lightweight convolutional neural networks such as EEGNet and ShallowNet  Schirrmeister et al., (2017) have achieved considerable performance in small EEG and MEG datasets.
+Using contrastive learning, it has been shown that merely using convolutional neural networks and projection layers can yield satisfactory results on neural datasets Chen et al., (2020).
+More recently, Benchetrit et al proposed a method towards real-time MEG-based reconstruction of visual perception Benchetrit Y, (2023).
 Song et al.
-presented an EEG encoder using ShallowNet spatiotemporal convolution module with a large convolution kernel with a few parameters for EEG embedding, resulting in favorable performance on EEG-based visual decoding [46].
+presented an EEG encoder using ShallowNet Temporal-Spatial convolution module with a large convolution kernel with a few parameters for EEG embedding, resulting in favorable performance on EEG-based visual decoding Song et al., (2023).
 Limitations of previous studies: Previous EEG studies are primarily oriented toward understanding visual perception in the human brain rather than maximizing EEG decoding performance.
 Thus the visual decoding performance is far from optimal.
-Specifically, previous studies have trained linear models to (1) classify a small set of images from brain activity [15, 23], (2) to predict brain activity from the latent representations of images [8], or (3) to quantify the similarity analysis between these two patterns with representational similarity [8, 15, 14, 3].
+Specifically, previous studies have trained linear models to (1) classify a small set of images from brain activity Grootswagers et al., (2019); King and Wyart, (2021), (2) to predict brain activity from the latent representations of images Cichy and Pantazis, (2017), or (3) to quantify the similarity analysis between these two patterns with representational similarity Cichy and Pantazis, (2017); Grootswagers et al., (2019); Gifford et al., (2022); Bankson et al., (2018).
 While these studies also utilize image embeddings, their linear decoders are limited to classifying a small group of object categories or distinguishing image pairs.
-Moreover, several deep neural networks have been applied to maximize classification of speech [10], cognitive load [21], and images [32, 29, 2] in EEG recordings.
-[32] proposed a deep convolutional neural network for classifying natural images using EEG signals.
-Unfortunately, the experiment presented all images of the same category in a single block, probably misleading the decoder to rely on autocorrelated noise rather than the hidden informative patterns of brain activity [27].
+Moreover, several deep neural networks have been applied to maximize classification of speech Défossez et al., (2023), cognitive load Jiao et al., (2018), and images Palazzo et al., (2020); McCartney et al., (2022); Bagchi and Bathula, (2022) in EEG recordings.
+Palazzo et al., (2020) proposed a deep convolutional neural network for classifying natural images using EEG signals.
+Unfortunately, the experiment presented all images of the same category in a single block, probably misleading the decoder to rely on autocorrelated noise rather than the hidden informative patterns of brain activity Li et al., (2020).
 Also, these EEG studies only classify a relatively small number of image categories.
 5 Discussion and Conclusion
-In this study, we developed an EEG-based image decoding and reconstruction framework.
-Employing self-supervised learning, our method enables the model to achieve better generalization performance in different tasks.
-To our knowledge, this is the first instance of EEG decoding achieving end-to-end image reconstruction with fMRI-level performance.
-Technical Impact: Our technical contributions are mainly on the EEG encoder and the two-stage image generator (Fig.
+In this study, we proposed a novel and feasible EEG-based zero-shot image reconstruction framework.
+Although it utilizes existing machine learning techniques , we demonstrate for the first time that EEG-based zero-shot visual decoding and reconstruction can be competitive with MEG and fMRI.
+Technical Impact: Our technical contributions are mainly on the EEG encoder and the two-stage zero-shot visual reconstruction framework (Fig.
 2).
-First, we developed the ATM, a plug-and-play EEG encoder which can efficiently extract EEG/MEG features for the three visual decoding tasks.
+First, we developed the ATM, an EEG encoder which can efficiently represent EEG/MEG features for three tasks.
 Our comprehensive experiments of the EEG encoder (Fig.
 3), compared to various architectures and training methods, achieves SOTA performance across various metrics and tasks (Figs.
-1b, 45).
-Second, the two-stage EEG guidance image generation achieves performance close to fMRI using only EEG data (Figs.
-67, Tab.
-3, 4), and this method is compatible with MEG data (Figs.
-1c, 8b).
-By comparing EEG embeddings and image embeddings obtained through prior diffusion, we demonstrated that EEG embeddings obtained via contrastive learning alone were insufficient for generating credible images (Appendix 0.C), suggesting that our two-stage generation strategy could effectively enhance generation performance even with limited data.
+4b, 5).
+Second, our two-stage EEG guidance image reconstruction framework achieves performance close to fMRI using only EEG data (Figs.
+614, Tab.
+1, 6), and this method is compatible with MEG data (Figs.
+4c, 7b).
 Neuroscience Insights: Our results offer insights into the relationship between brain activity and visual perception.
 We analyzed EEG-based visual decoding within different time windows to examine when visual information is perceived in the brain (Fig.
+7).
+Our results revealed that visual information in EEG data is predominantly contained within the 200-400ms range (Fig.
+7a), consistent with previous EEG studies Teichmann et al., (2023); Grootswagers et al., (2022); Song et al., (2023).
+Interestingly, the visual information in MEG data last up to 800ms, much longer than EEG (Fig.
+7b), in line with the results reported by a previous MEG study Benchetrit Y, (2023); Song et al., (2023).
+We also found that EEG performs better than MEG in visual tasks (See Appendix D for Tab.
+6), which is different from other fields, such as speech decoding Défossez et al., (2023).
+In addition, through ablation experiments of spatial information, we found that visual information is mainly encoded in the occipital and parietal areas (Fig.
 8).
-Our results revealed that visual information in EEG data is predominantly contained within the 200-400ms range (Fig.
-8a), consistent with previous EEG studies [49, 16, 46].
-Interestingly, the visual information in MEG data last up to 800ms, much longer than EEG (Fig.
-8b), in line with the results reported by a previous MEG study [4, 46].
-We also found that EEG performs better than MEG in visual tasks (See Appendix 0.D for Tab.
-4), which is different from other fields, such as speech decoding [10].
-In addition, through ablation experiments of spatial information, we found that visual information is mainly encoded in the occipital and parietal areas (Fig.
-9).
-Using the image reconstruction tasks, we can visualize the information encoded in the brain, providing a window for vision researchers to explore how humans perceive visual information.
-Interesting Phenomena and Future Directions: Our study uncovered intriguing phenomena, sheding lights on the future direction of EEG decoding.
-First, there are non-negligible performance differences between cross-subject and within-subject settings.
-This performance gap arises from inter-subject differences in EEG signals [13, 58], likely attribute to heterogeneity in individual brain, differences in visual perception between individuals, and even shifts in noise distribution during EEG recording.
-To address this cross-subject challenge, it calls for more efforts on EEG encoder, such as more flexible neural network architectures or training with larger EEG dataset.
-Transfer learning and meta-learning are also future directions worth exploring [55, 54, 57].
-On the other hand, large EEG models pre-trained on massive EEG datasets may be the ultimate solution for visual decoding tasks [9, 7].
-Nevertheless, how to unify various electrode montages of different EEG datasets when pre-training large EEG models is a challenge.
-EEG source localization, which converts senor-level EEG signals into the standard brain source space [52, 50], might be a potential solution.
-Moreover, our EEG-based image reconstruction results are somewhat limited by the visual features used from CLIP.
-For instance, it is easy to reconstruct semantics but difficult to preserve low-level visual feature aspects.
-This limitation leads to generated images lacking detail and richness.
-Future studies could use multiple different visual features to align EEG and image data.
-This proposition finds support in similar methodologies applied in fMRI-based studies [40, 26], which combine high-level semantic features extracted by CLIP with low-level visual features from VAE to improve consistency.
-Acknowledgements
-This work is supported by the National Key R&D Program of China (2021YFF1200804),
-Shenzhen Science and Technology Innovation Committee (20200925155957004,
-KCXFZ20201221173400001, SGDX2020110309280100).
-This work originated as a course project in the BI&AI 2023 course at SUSTech, and the author would like to thank course project members, Yuanhao Fan and Zhihong Wu, for their contributions to the early stages of this work.
-Thank Haoyang Qin for his time and advice.
-Appendix 0.A Datasets for experiments
-0.A.1 EEG dataset
-We conducted our experiments on the THINGS-EEG dataset’s training set [14, 16].
+Interesting Phenomena and Future Directions: First, there are non-negligible performance differences between cross-subject and within-subject settings.
+This performance gap arises from inter-subject differences in EEG signals Gibson et al., (2022); Xu et al., (2020), likely attribute to heterogeneity in individual brain, differences in visual perception between individuals, and even shifts in noise distribution during EEG recording.
+So it calls for more efforts on EEG encoder, such as more flexible neural network architectures or better weight initialization of pre-trained models Cui et al., (2023); Chen et al., (2024).
+Transfer learning and meta-learning are also future directions worth exploring Wu et al., (2020, 2022); Xie et al., (2023).
+Moreover, how to unify various electrode montages of different EEG datasets when pre-training large EEG models is a challenge.
+EEG source localization, which converts senor-level EEG signals into the standard brain source space Wei et al., (2021); Wang et al., (2024), might be a potential solution.
+Supplementary Material:
+Visual Decoding and Reconstruction via EEG Embeddings with Guided Diffusion
+Appendix A Datasets for experiments
+A.1 EEG dataset
+We conducted our experiments on the THINGS-EEG dataset’s training set Gifford et al., (2022); Grootswagers et al., (2022).
 This dataset includes a large EEG corpus from 10 human subjects during the visual task.
-The experiment employed the Rapid Serial Visual Presentation (RSVP) paradigm for orthogonal target detection tasks to ensure that participants attended to the visual stimuli.
+The experiment employed the Rapid Serial Visual Presentation (RSVP) paradigm for orthogonal target detection tasks to ensure that participants attended to the visual stimulus.
 All 10 participants completed 4 equivalent experiments, resulting in 10 datasets with 16,540 training image conditions repeated 4 times, and 200 testing image conditions repeated 80 times, totaling (16,540 training image conditions $\times$ 4 repetitions) + (200 testing image conditions $\times$ 80 repetitions) = 82,160 image trials.
@@
 Baseline correction was performed using the mean of the 200 ms pre-stimulus data.
-All electrodes were retained and downsampled to 250 Hz for analysis, and multivariate noise normalization was applied to the training data [17].
-To improve signal-to-noise ratio, we averaged across the four EEG trials from the same image in the test set, while keeping each EEG trial in the training setting.
+All electrodes were retained and downsampled to 250 Hz for analysis, and multivariate noise normalization was applied to the training data Guggenmos et al., (2018).
+To improve signal-to-noise ratio, we averaged across the 4 EEG trials from the same image in the test set, while keeping each EEG trial in the training setting.
 We compared the effects of averaging across EEG trials and found it indeed improved the performance.
-0.A.2 MEG dataset
-To verify the versatility of ATM for embedding electrophysiological data, we tested it on MEG data modality using the THINGS-MEG dataset [18].
+A.2 MEG dataset
+To verify the versatility of ATM for embedding electrophysiological data, we tested it on MEG data modality using the THINGS-MEG dataset Hebart et al., (2023).
 It includes 271-channel MEG data from 4 subjects with 12 MEG sessions.
@@
 We compared our approach with advanced methods i.e.
-NICE [46] and B.D.
-[4] for classification and retrieval tasks on the MEG dataset.
-Similar to [46], we directly used the stimulus images to match the template, rather than other images belonging to the concept.
-Appendix 0.B More Implementation Details
-0.B.1 Evaluation metric implementation
+NICE Song et al., (2023) and B.D.
+Benchetrit Y, (2023) for classification and retrieval tasks on the MEG dataset.
+We directly used the stimulus images to match the template, rather than other images belonging to the concept.
+Appendix B More Implementation Details
+B.1 Evaluation metric implementation
 Classification accuracy
@@
 Specifically, we assessed performance based on the Top-k (where k=1, 5) predictions.
-This means a classification is considered correct if the true category is among the model’s Top-k predicted categories.
 We conducted tests for both within-subject and leave-one-subject-out classification accuracy, enabling a comprehensive evaluation of the model’s performance across different scenarios.
@@
 This means, apart from the entire test set, the model evaluated by N-Way accuracy (where N=2, 4, 10 in our experiments) on the test set.
-We report these results in Appendix 0.G.
+We report these results in Appendix H.
 Retrieval accuracy
 Similar to the classification task, in the retrieval task, the objective is to retrieve the Top-K images most related to a given stimulus image via its corresponding EEG signal.
-This implies that by merely changing the text embeddings of image labels used during prediction, to image embeddings, we can transition the task from classification to image retrieval.
-The performance in retrieval tasks is superior to classification tasks as using image embeddings in training.
-We conducted a detailed comparison of different methods in terms of their Top-K retrieval accuracy in image retrieval tasks.
-Given that contrastive learning is known to be sensitive to batch size, we also compared the performance improvement of different methods under varying batch sizes (batch size=16, 1024) (Appendix 0.G).
+This implies that by changing the text embeddings of image labels to image embeddings, we can transition the task from classification to image retrieval.
+Given that contrastive learning is known to be sensitive to batch size, we also compared the performance improvement of different methods under varying batch sizes (batch size=16, 1024) (Appendix H).
 Generation accuracy
 The generation task presents more challenges than the other tasks.
-In this phase, EEG representations are fed into a two-stage generator.
-For each image condition in the test set, we generate 1 images from 10 subjects based on the corresponding EEG signal.
+For each image condition in the test set, we generate 10 different images from 10 subjects based on the corresponding EEG.
 Subsequently, image retrieval is performed for each generated image.
 The Top-1 and Top-5 accuracies are calculated.
-Since we utilized CLIP for alignment and did not extract lower-level visual features from the EEG, this metric is particularly useful.
 It helps in evaluating the semantic alignment between the generated images and their original counterparts.
-0.B.2 Computing methods implementation
+B.2 Computing methods implementation
 In the upstream EEG encoder part, we compared various methods.
 For the B.D.
-method  [4], we replicated the network structure as described in the original work, with the difference being in the shape of the input data, due to the original study’s focus on MEG.
-It is worth mentioning that we used the retention test method in the testing process, so the strategy of joint subject training is not suitable for our task.
-We modify its subject-wise layer as an Input layer for modeling the time dimension.
-Regarding the  [46], we also reproduced the EEG encoder as described in their paper, using spatiotemporal convolution modules and EEG projection modules with the same convolution parameters.
+method  Benchetrit Y, (2023), we replicated the network structure as described in the original work, with the difference being in the shape of the input data due to the original study’s focus on MEG.
+It is worth mentioning that we used the leave one for subject method in the testing process so we modify its subject-wise layer as an linear layer for modeling the time dimension.
 To ensure fairness, we did not use the same hyperparameters as in the original paper.
-Instead, we chose settings that yielded excellent results upon reproduction.
+Instead, we chose settings yielded excellent results upon reproduction.
 Across all methods, we used identical hyperparameters, apart from the network structures.
 These included batch size, optimizer, initial learning rate, and temperature parameters.
-0.B.3 Architecture details
-Table 1: Brain module configuration for use with a target latent of size 1024
+B.3 Architecture details
+Table 2: Brain module configuration
 Layer Input shape Output shape # parameters
-Spatial attention block (N, C, T) (N, C, T) 553,078
-Input layer (N, C, T) (N, C, T) 62,750
-Plug-and-Play Conv module (N, C, T) (N, H1, H2) 103,680
-Temporal aggregation (N, H1, H2) (N, H1*H2) 0
+Channel-wise attention layer (N, C, T) (N, C, D) 553,078
+Temporal-Spatial Conv module (N, C, D) (N, H1, H2) 103,680
+Temporal-Spatial aggregation (N, H1, H2) (N, H1*H2) 0
 MLP projector (N, H1*H2) (N, 1024) 2,527,232
-Total 3,246,741
+Total 3,183,990
 Table 3: Ablation study on the ATM model’s different components for THINGS-EEG retrieval.
-Module MLP AdaConv IL SA TOP-1 TOP-5
-8.01 ±1.97 25.41±5.74
-21.55±6.26 50.78±8.77
-23.57±6.00 53.50±7.26
-27.71±6.22 58.71±7.49
-ATM-S 26.49±5.81 56.58±8.67
-8.01 ±1.97 25.41±5.74
-18.67±5.48 46.31±9.11
-21.11±5.56 50.48±7.74
-24.64±5.97 54.17±5.93
-ATM-E 22.43±5.41 51.24±7.27
-0.B.4 Training details
-In the initial design of the EEG encoder module, we adopted two approaches to guide model predictions: text embedding and image embedding.
-Due to the differences in feature granularity, alignments favoring image embedding tend to perform better in image retrieval and classification tasks.
-During the training process using contrastive learning, we found that a batch size of 16 is a prudent choice across all models.
-A batch size of 1024 means that a sufficient number of samples are compared in a single training step, which requires the model to have a higher noise resistance capability.
-As observed in Tab.
-5 and Tab.
-6, larger batch sizes have the potential to achieve better performance and definite training efficiency gains across most methods.
-To improve the signal-to-noise ratio of EEG, we averaged 80 repeated data in the test set, a method similar to seeking Event-Related Potentials (ERP).
-To make full use of the training data volume, we did not average the 4 repetitions in the training set but instead fed all EEG data into the model for learning.
-Figure 10: Test accuracy during training.
-(a) Training of within-subject model.
-(b) Training of across-subject model.
-We compared 5 different EEG encoding models, including EEGNetv4, B.D., NICE, ATM-S and ATM-E.
-Appendix 0.C Details of EEG guidance image generation
-Here, we provide a concise overview of the conditional diffusion model framework used in EEG-guided image generation, following the presentation of continuous-time diffusion models in [45, 22].
+Module MLP TSConv CAL TOP-1 TOP-5
+8.11±1.74 26.83±4.78
+21.65±6.22 51.34±9.83
+ATM-S 23.73±7.62 52.71 ±9.71
+28.64±6.39 58.47±8.97
+8.11±1.74 26.83±4.78
+17.95±5.95 43.10±8.63
+ATM-E 23.73±7.62 52.71 ±9.71
+24.92±6.00 54.78±8.49
+B.4 Model configuration
+To validate the efficacy of our EEG encoder, we experimented with a variety of empirical setups aimed at optimizing the model’s efficiency.
+we leverage joint subject training to adapt to new subjects.
+Once a model is trained, it can be used for both reasoning about known subjects (subject-specific tokens) and reasoning about unknown subjects (shared tokens).
+In the context of channel-wise attention layer, we explored four distinct approaches for enhancing downstream retrieval capabilities: leveraging subject-specific tokens for retrieval, averaging all tokens for a more generalized retrieval, flattening the entire token set for direct retrieval, and preserving token dimensions to feed into the Temporal-Spatial convolution module for feature integration.
+Notably, the strategy of preserving dimensions and using the Temporal-Spatial convolution module emerged as the most effective.
+In our quest to optimize token embeddings, we experimented with a variety of approaches, including using $1\times 1$ convolution and linear layers.
+Under our framework, using linear layers performs better than convolutions.
+Moving beyond, we delved into the efficacy of diverse Feed Forward Networks (FFNs) within the Transformer encoder layer.
+Our findings indicated that an FFN tailored to the temporal dimension emerged as the superior option.
+We also conducted a comparative analysis of various positional encoding techniques.
+Interestingly, for Temporal-Spatial convolution utilized in retrieval tasks, the significance of positional encoding diminished.
+In a further exploration, we examined the impact of deploying convolutions at various stages and even contemplated the complete removal of the convolution module.
+It was discovered that situating convolutions post the Transformer encoder layer yielded the most favorable outcomes.
+Conversely, a shift to a MLP or the removal of the convolution module led to a notable degradation in performance.
+Our assumption is though the convolution’s inherent translational invariance is compromised when the context of the time dimension is disrupted, the efficiency of parameters it possesses may confer a resistance to overfitting, thereby maintaining its effectiveness.
+Table 4: Impact of each module on the result in different configurations.
+Module Config Top-1 (std) Top-5 (std)
+Channel-wise attention layer w/ mean token 7.29 (3.01) 23.11 (6.62)
+w/ flatten token 14.85 (5.46) 37.56 (8.92)
+w/ keep dim 28.64 (6.39) 58.47 (8.97)
+Token embedding w/ conv1d 24.81 (7.29) 55.68 (9.08)
+w/ linear 28.64 (6.39) 58.47 (8.97)
+Feed Forward Network w/ temporal dim 28.64 (6.39) 58.47 (8.97)
+w/ spatial dim 19.92 (7.74) 47.59 (11.5)
+w/o 26.45 (7.73) 57.00 (9.95)
+Position encoding w/ sinusoidal 27.96 (6.54) 58.16 (8.44)
+w/ learnable absolute 26.66 (6.56) 56.95 (7.95)
+w/o 28.64 (6.39) 58.47 (8.97)
+Temporal spatial convolution w/ pre 14.23 (4.20) 37.44 (6.06)
+w/ post 28.64 (6.39) 58.47 (8.97)
+w/ both 25.37 (5.16) 57.07 (6.13)
+w/o 23.73 (7.62) 52.71 (9.71)
+B.5 Training details
+In our EEG projector module, we integrated two distinct strategies for steering model predictions: text embedding and image embedding.
+Given the variance in feature granularity, we observed that alignments that prioritize image embedding excel in tasks of image retrieval and classification.
+Throughout the training phase, our experiments revealed that a batch size of 16 is a judicious selection for all models.
+Conversely, a batch size of 1024, which implies a substantial number of samples are processed in each training iteration, necessitates the model to exhibit a heightened capacity for noise resistance.
+In order to enhance the signal-to-noise ratio within EEG data, we implemented an averaging technique on 80 repeated instances within the test set.
+This approach mirrors the methodology employed in identifying Event-Related Potentials (ERP).
+To maximize the utilization of the available training data, we refrained from averaging the 4 repetitions in the training set.
+Instead, we opted to input the complete set of EEG data into the model, thereby facilitating comprehensive learning.
+Figure 9: Test accuracy during each training epoch.
+(a) Training of different within-subject models.
+(b) Training of different across-subject models.
+We compared 7 different EEG encoding models, including EEGconformer, MLP, EEGNetv4, B.D., NICE, ATM-S (Ours) and ATM-E (Ours).
+Appendix C Details of EEG guidance image generation
+Here, we provide a concise overview of the conditional diffusion model framework used in EEG-guided image generation, following the presentation of continuous-time diffusion models in Song et al., (2020); Karras et al., (2022).
 Diffusion models
@@
 Training of DMs
-The core of DM training is to learn a model $s_{\theta}(x;t)$ for the score function.
+The core of DM training is to learn a model for the score function.
 This is typically achieved through denoising score matching (DSM), where $\epsilon_{\theta}$ is a learnable denoiser.
 The training process can be formulated as:
-$\mathbb{E}_{(x_{0},c)\sim p_{\text{data}}(x_{0},c),(n_{t},t)\sim p(n_{t},t)}\left[\|\epsilon_{\theta}(x_{0}+n_{t};t,c)-x_{0}\|_{2}^{2}\right],$ (3)
-where $n_{t}$ is Gaussian noise with variance $\sigma_{t}^{2}$, and $c$ represents a condition.
-0.C.1 Stage I - EEG-Conditioned Diffusion
+$\mathbb{E}_{(x_{0},c)\sim p_{\text{data}}(x_{0},c),(n_{t},t)\sim p(n_{t},t)}\left[\|\epsilon_{\theta}(x_{0}+\sigma_{t}\epsilon;t,c)-\epsilon\|_{2}^{2}\right],$ (3)
+where $\epsilon$ is Gaussian noise with variance $\sigma_{t}^{2}$, and $c$ represents a condition.
+C.1 Stage I - EEG-Conditioned Diffusion
 The initiation of the EEG-conditioned diffusion phase is paramount in our EEG-based image generation framework, leveraging the classifier-free guidance strategy alongside data pairs of CLIP embeddings and EEG embeddings $(z_{I},z_{E})$.
@@
 The primary objective of this method is to enhance the sample quality produced by DMs while maintaining output diversity.
-0.C.2 Stage II - CLIP-Embedded Image Synthesis
+C.2 Stage II - CLIP-Embedded Image Synthesis
 In Fig.
-11, we compare the effects of one-stage and two-stage EEG-guided image generation.
+10, we compare the effects of one-stage and two-stage EEG-guided image generation.
 We show images generated using EEG embeddings directly (One-stage) and images generated using image embeddings obtained via prior diffusion (Two-stage).
 It can be seen that the two-stage EEG-guided image generation can more accurately reconstruct the semantic and low-level visual features of the original image, and the style is more realistic.
-Figure 11: Comparison between one-stage and two-stage EEG guidance image generation.
-(a) We present the images that subjects seen (Seen), the generated images directly using EEG embeddings (One-stage), and the generated images from image embeddings obtained by the prior diffusion (Two-stage).
-These results indicate that the strategy of our two-stage generation can better reconstruct the seen visual stimuli.
+Figure 10: Comparison between one-stage and two-stage EEG guidance image reconstruction.
+(a) We present the images that subjects seen (Seen), our reconstructed images directly using EEG embeddings (One-stage), and the reconstructed images from low level and high level image embeddings obtained by the prior diffusion (Two-stage).
+These results indicate that the strategy of our two-stage generation can better reconstruct the seen visual stimulus.
 (b) We employed ATM-S to compare the generated images with the original images in a retrieval task.
 Our result indicates that the images generated in two stages significantly enhance the performance of the original model on the retrieval task.
+Table 5: Latent VAE retrieval performance
+Condition Top-1 (%) Top-5 (%)
+ideal chance 0.5 2.5
+ATM-S (Ours) 10.14 29.55
 In the second stage of our EEG-based image generation approach, the CLIP embedding $z_{I}$ derived from the EEG-conditioned diffusion acts as the precursor for synthesizing visual objects $I$ based on $z_{I}$.
-This is achieved by harnessing the synergies of advanced pre-trained models, namely SDXL and IP-Adapter [33, 59], facilitating the creation of high-caliber images.
+This is achieved by harnessing the synergies of advanced pre-trained models, namely SDXL and IP-Adapter Podell et al., (2023); Ye et al., (2023), facilitating the creation of high-caliber images.
 The cornerstone of our synthesis process is the SDXL framework, acclaimed for its proficiency in text-to-image conversion.
@@
 SDXL-turbo for accelerated processing
-To augment the efficiency of our framework, we additionally explore the SDXL-Turbo [36], a refined iteration of SDXL optimized for swift image synthesis.
+To augment the efficiency of our framework, we additionally explore the SDXL-Turbo Sauer et al., (2023), a refined iteration of SDXL optimized for swift image synthesis.
 This variant proves especially beneficial in scenarios demanding quick generation of high-fidelity visuals.
@@
 Its compatibility with text prompts for multimodal image generation extends the versatility of our EEG-based image synthesis approach.
-Appendix 0.D Performance comparison
+C.3 Low-level pipeline
+Compared with pure vision pre-training models such as (ViT, ResNet, DINO, etc.), the CLIP model lacks low-level visual features.
+Therefore, in order to make up for this shortcoming, our framework introduces a low-level visual reconstruction pipeline.
+We hope to restore basic such as contour, posture, orientation and other pixel-level information from EEG by aligning with the latent of VAE.
+Past work Takagi and Nishimoto, (2023) has found that in the denoising stage of the early diffusion model, $z$ signals (corresponding to the VAE latent in our framework) dominated prediction of fMRI signals.
+And during the middle step of the denoising process, $z_{c}$ predicted activity within higher visual cortex much better than $z$.
+However, note that this is only an analysis based on decoding accuracy.
+These analyses do not have a strong neuroscience causal relationship.
+We still cannot conclude that the low-level features of neural data are modeled by VAE.
+We trained the low-level pipe for 200 epochs, trying a latent mean squared error (MSE) loss, along with a contrastive learning loss, and a variational autoencoder (VAE) image reconstruction loss to align the $4\times 64\times 64$ EEG latents obtained from a projection layer and an upsampled CNN with the VAE latents.
+Nevertheless, reconstruction loss or contrastive learning loss performs worse than only applying the loss in latent space and also requires significantly more GPU memory.
+In addition, we found that using a low-level visual model for distillation learning in the low-level pipeline is not only unhelpful for VAE latent training, but also leads to overfitting.
+Similar conclusions were reached in MindEyeScotti et al., (2023).
+Our results suggested that the low-level zeroshot reconstruction in EEG is not stable enough and may mislead the model results.
+When using the low level pipeline, we usually set the inference steps of SDXL to $10$ (or SDXL-turbo to $4$) and the image-to-image denoising strength to $0.5$.
+We give several reconstruction examples in Fig.
+11 to compare the impact before and after using low level pipeline.
+Moreover, low level alignment was validated through retrieval performance tests, as depicted in Tab.
+5.
+Our findings indicate that retrieval using EEG latents can also achieve excellent performance.
+This further substantiates the feasibility of aligning the low-level consistency through VAE latents.
+Figure 11: Example of our reconstructions for Subject 8 output from different pipelines.
+From left to right: reconstruction using only CLIP (i.e., Only CLIP), using only CLIP and the semantic pipeline (i.e., CLIP $+$ Semantic), using only the low level pipeline (i.e., Only low level), using only CLIP and the low level pipeline (i.e., CLIP $+$ low level), using joint CLIP, low level, and semantic pipelines.
+C.4 Semantic-level pipeline
+In addition to using EEG latent and low-level pipelines during reconstruction, we also add a corresponding semantic level pipeline guided by text captions during the image reconstruction.
+We input the $1\times 1024$ EEG features output by prior Diffusion into the trained image projector to obtain $256\times 1024$ image features.
+Using the GIT modelWang et al., (2022), we can directly generate a caption from the latent features of the image.
+IP-Adapter accepts such a caption as a text prompt to guide the semantic level reconstruction of the image.
+It should be noted that due to the difficulty of the zeroshot task itself and the low dimensionality of the EEG features, the caption generated from the latent may be unstable, thereby interfering with the original correct EEG semantics.
+Considering that the image features extracted by the CLIP model itself are already high-level visual features and do not require the introduction of more semantic information, this framework retains the entry of the text prompt, and the reconstructed image presented does not force the use of the semantic level pipeline.
+Appendix D Performance comparison
 Comparison metrics
-Our study uses various metrics to evaluate how well we can recreate visual stimuli from brain data (EEG, MEG, fMRI) (Table 1 in main text).
-These metrics include SSIM (structural similarity index metric)[51], SwAV (SwAV-ResNet50, refer to average correlation distance)[5], and two-way identification using neural networks (AlexNet(2/5), Inception, CLIP.
+Our study uses various metrics to evaluate how well we can recreate visual stimulus from brain data (EEG, MEG, fMRI) (Tab.
+1 and Tab.
+7).
+These metrics include PixCorr (pixelwise correlation, between ground truth and reconstructions), SSIM (structural similarity index metric)Wang et al., (2004), SwAV (SwAV-ResNet50, refer to average correlation distance)Caron et al., (2020), and two-way identification using neural networks (AlexNet(2/5), Inception, CLIP.
 Here AlexNet(2/5) the 2nd and 5th feature layers of AlexNet) for both low-level and high-level image features.
-Here two-way identification can be seem as a two-way retrieval task described in Appendix  0.B.1.
+Here two-way identification can be seem as a two-way retrieval task described in Ozcelik and VanRullen, (2023).
 In Tab.
-3, our results showed that on the THINGS dataset, we could achieve performance over MEG on EEG reconstruction using ATM.
+1, our results showed that on the THINGS dataset, we could achieve performance over MEG on EEG reconstruction using ATM.
 Tab.
 6 shows the decoding performance of different data sets (fMRI, MEG, EEG) on visual stimulus tasks, and we even achieved the same or better performance than fMRI and MEG.
-Our results suggest that a suitable upstream neural representation plays a decisive role in the downstream task.
-Table 3: Quantitative assessments of the reconstruction quality for EEG, MEG, and fMRI. For detailed explanations of the metrics.
-Low-level High-level
-Dataset $\uparrow$ SSIM $\uparrow$ AlexNet(2) $\uparrow$ AlexNet(5) $\uparrow$ Inception $\uparrow$ CLIP $\uparrow$ SwAV $\downarrow$
-NSD (B.D.) [4] 0.366 0.962 0.977 0.910 0.917 0.410
-NSD (Brain-Diffuser) [31] 0.356 0.942 0.962 0.872 0.915 0.423
-NSD (MindEye) [39] 0.308 0.917 0.974 0.936 0.942 0.369
-THINGS-MEG (B.D.) [4] 0.327 0.695 0.753 0.593 0.700 0.630
-THINGS-MEG (Ours) 0.340 0.613 0.672 0.619 0.603 0.651
-THINGS-EEG (Ours) 0.345 0.776 0.866 0.734 0.786 0.582
+Our results suggest that a suitable neural representation plays a decisive role in the downstream task.
 Table 6: The classification performance of various methods are discussed.
@@
 Dataset Model top-1 top-5 top-1 top-5 top-1 top-5
-GOD-Wiki (fMRI) CADA-VAE (V&T)[38] 10.02 40.37 - - - -
-MVAE (V&T) [56] 10.04 39.60 - - - -
-MMVAE (V&T) [43] 11.68 43.29 - - - -
-MoPoE-VAE (V&T) [47] 12.90 51.78 - - - -
-BraVL (V&T) [11] 13.99 53.13 - - - -
+GOD-Wiki (fMRI) CADA-VAE (V&T)Schonfeld et al., (2019) 10.02 40.37 - - - -
+MVAE (V&T) Wu and Goodman, (2018) 10.04 39.60 - - - -
+MMVAE (V&T) Shi et al., (2019) 11.68 43.29 - - - -
+MoPoE-VAE (V&T) Sutter et al., (2020) 12.90 51.78 - - - -
+BraVL (V&T) Du et al., (2023) 13.99 53.13 - - - -
 THINGS (MEG) ATM (Ours) 15.63 41.38 11.75 29.25 5.88 19.25
-THINGS (EEG) BraVL [11] 14.33 40.28 - - 5.82 17.45
+THINGS (EEG) BraVL Du et al., (2023) 14.33 40.28 - - 5.82 17.45
 ATM (Ours) 17.40 39.40 11.50 28.50 7.40 20.60
-Appendix 0.E Representational analysis
+Table 7: Quantitative comparison results of image reconstruction in Subject 8 via our framework using different encoders.
+Low-level High-level
+Dataset $\uparrow$ PixCorr $\uparrow$ SSIM $\uparrow$ AlexNet(2) $\uparrow$ AlexNet(5) $\uparrow$ Inception $\uparrow$ CLIP $\uparrow$ SwAV $\downarrow$
+NSD-fMRI Benchetrit Y, (2023) 0.305 0.366 0.962 0.977 0.910 0.917 0.410
+THINGS-MEG Benchetrit Y, (2023) 0.058 0.327 0.695 0.753 0.593 0.700 0.630
+THINGS-MEG (averaged) Benchetrit Y, (2023) 0.090 0.336 0.736 0.826 0.671 0.767 0.584
+THINGS-MEG (Ours) 0.104 0.340 0.613 0.672 0.619 0.603 0.651
+[…]
+THINGS-EEG (EEGNetV4)Lawhern et al., (2018) 0.140 0.302 0.767 0.840 0.713 0.773 0.581
+THINGS-EEG (Ours) 0.160 0.345 0.776 0.866 0.734 0.786 0.582
+Appendix E Representational analysis
 As depicted in Fig.
@@
 (b) Visualization in the latent space of EEG/image/text by t-SNE.
-Appendix 0.F Additional images results
-0.F.1 Additional retrieval results
+Appendix F Concept analysis
+We have adopted the concept embedding encoder proposed by Wei et al.
+Wei et al., (2024), which encodes the clip embedding of the original image into a 42-dimensional vector, with each dimension representing a distinct concept.
+Utilizing the ATM for direct projection on the EEG data of 200 categories from the test set, we obtained EEG embeddings that were then fed into the concept encoder with frozen weights, yielding 200 concept embeddings, as depicted in Fig.
+13.
+An analysis of representational similarity at the concept level indicates that the EEG embeddings derived from our EEG projector effectively align with the conceptual space.
+This ensures semantic consistency at a high level of alignment, providing compelling evidence for the reconstruction of images from EEG data.
+Figure 13: Visualization of the conceptual representation analysis.
+(a) Conceptual representations were obtained from eeg embeddings using concept encoder.
+(b) The similar matrix between EEG embeddings and real concept embeddings.
+(c) Concept embedding similarity matrix after cluster rearrangement (k=5).
+Appendix G Additional images results
+We visualize the best, medium and worst generated images in Fig.
+14.
+We randomly selected the EEG data of a subject viewing 100 images, and extracted EEG embeddings to guide image generation.
+By calculating the cosine similarity of the CLIP embedding between the generated image and the original image, we found 12 images each with the best, medium and worst generation effects.
+It can be seen that in the best group, the generated image is not only highly consistent with the semantics of the original image, but also well retains the low-level visual features.
+in the medium group, the generated image maintains the semantic features of the original image, and the low-level visual features are well preserved.
+Visual features were altered.
+in the worst group, both semantic features and low-level visual features were altered.
+Figure 14: Examples of EEG-guided visual reconstruction.
+From top to bottom, we exhibit the best, median, and worst 12 generated images, respectively.
+We show the images subjects seen and the generated images by our two-stage image generator.
+G.1 Additional retrieval results
 Figure 15: Additional retrieval results
-0.F.2 Additional generated images
+G.2 Additional generated images
 Figure 16: Additional generated results with the best alignment to original images
@@
 Figure 18: Additional generated results with the worst alignment to original images
-0.F.3 Additional generated images for each subject
+G.3 Additional generated images for each subject
 Figure 19: Part of subject 1 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
@@
 Figure 28: Part of subject 10 generates images. We do a batch generation of the subjects and then calculate the best, medium, and worst performers compared to the original stimulus pictures.
-Appendix 0.G Additional evaluation results
-0.G.1 Accuracy for time windows
+Appendix H Additional evaluation results
+H.1 Accuracy for time windows
+According to our results in 30 for time windows, which shows that for all embeddings, a clear peak can be observed for windows ending around 200-250 ms after image onset.
+Comparing 29 and 30, we can see that, unlike Benchetrit Y, (2023), our time window results on THINGS-EEG dataset do not have a second peak, which may be mainly affected by the experimental paradigm.
+In 7, we can see that in the first 50ms-200ms, the image reconstructed at 50 ms is completely messy and has no semantics; the semantics of the reconstructed image after 250 ms is basically correct and gradually stabilizes, and after 500ms, due to the lack of additional visual response, the content of the image reconstruction is more stable.
+Similar results are also shown in Figure 4 A of Benchetrit Y, (2023): their method can also reconstruct high-quality images at 100ms.
+This just shows that our reconstruction results are in line with the neuroscience prior.
+However, this does not mean that the EEG data after the absence of visual response (200ms) loses its contribution to decoding, because the processing of high-level visual features (corresponding to the visual features of CLIP) may be involved over time.
 Figure 29: Accuracy for growing windows. We use an EEG time window of 100ms, sliding 100ms each time. (a) Top-1 accuracy. (b) Top-5 accuracy.
@@
 Figure 32: Accuracy for sliding windows. We use an MEG time window of 100ms, sliding 100ms each time. (a) Top-1 accuracy. (b) Top-5 accuracy.
-Table 5: Overall accuracy of zero-shot retrieval on THINGS-EEG dataset.
-We showed in-subject and cross-subject retrieval task performance (Ave $\pm$ Std.%) under the condition of batch size=16.
-We compared the 2-way, 4-way, 10-way, the Top-1 and Top-5 accuracy of 200-way from different EEG embedding methods.
-Our ATM outperformed all the others.
-Subject dependent - train and test on one subject (batch size=16)
-Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet
-83.23 $\pm$ 2.77
-65.67 $\pm$ 3.84
-43.87 $\pm$ 3.78
-7.79 $\pm$ 1.27
-22.12 $\pm$ 2.88
-EEGConformer
-89.18 $\pm$ 2.88
-75.49 $\pm$ 5.30
-55.32 $\pm$ 6.88
-11.29 $\pm$ 3.55
-33.49 $\pm$ 6.78
-ShallowFBCSPNet
-84.29 $\pm$ 2.56
-67.13 $\pm$ 3.61
-46.07 $\pm$ 4.55
-8.26 $\pm$ 2.95
-25.87 $\pm$ 5.03
-EEGNetV4
-89.03 $\pm$ 2.86
-75.40 $\pm$ 5.18
-56.77 $\pm$ 6.60
-13.29 $\pm$ 3.99
-35.50 $\pm$ 7.14
-B.D.
-91.14 $\pm$ 2.45
-79.59 $\pm$ 4.59
-62.62 $\pm$ 6.33
-16.29 $\pm$ 4.35
-42.16 $\pm$ 7.56
-NICE
-92.17 $\pm$ 2.78
-82.10 $\pm$ 5.55
-65.65 $\pm$ 8.03
-19.40 $\pm$ 5.83
-46.26 $\pm$ 10.42
-MLP
-87.67 $\pm$ 4.41
-72.58 $\pm$ 7.62
-53.52 $\pm$ 9.45
-11.81 $\pm$ 4.33
-32.74 $\pm$ 9.24
-ATM-S (Ours)
-93.89 $\pm$ 1.97
-85.38 $\pm$ 3.95
-71.63 $\pm$ 6.17
-22.84 $\pm$ 5.87
-52.22 $\pm$ 8.34
-ATM-E (Ours) 94.44 $\pm$ 1.76 86.11 $\pm$ 3.39 71.89 $\pm$ 5.64 24.38 $\pm$ 6.56 54.14 $\pm$ 8.08
-Subject independent - leave one subject for test (batch size=16)
-Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet
-77.14 $\pm$ 4.14
-55.29 $\pm$ 6.22
-31.88 $\pm$ 6.20
-2.93 $\pm$ 1.70
-13.00 $\pm$ 4.45
-EEGConformer
-79.49 $\pm$ 3.91
-59.31 $\pm$ 5.83
-36.67 $\pm$ 5.35
-4.11 $\pm$ 1.83
-17.04 $\pm$ 4.21
-ShallowFBCSPNet
-75.14 $\pm$ 4.45
-53.06 $\pm$ 6.70
-31.23 $\pm$ 6.56
-3.19 $\pm$ 2.17
-13.01 $\pm$ 5.15
-EEGNetV4
-82.60 $\pm$ 3.17
-64.28 $\pm$ 5.44
-42.24 $\pm$ 6.10
-6.13 $\pm$ 2.40
-21.23 $\pm$ 5.19
-B.D.
-81.49 $\pm$ 3.52
-62.35 $\pm$ 6.44
-40.53 $\pm$ 7.19
-6.16 $\pm$ 2.40
-20.45 $\pm$ 5.40
-NICE
-81.85 $\pm$ 2.53
-63.57 $\pm$ 4.46
-41.86 $\pm$ 4.68
-6.43 $\pm$ 1.46
-21.39 $\pm$ 3.62
-MLP
-80.49 $\pm$ 2.61
-62.09 $\pm$ 3.88
-40.37 $\pm$ 3.71
-5.67 $\pm$ 1.09
-19.97 $\pm$ 3.02
-ATM-S (Ours)
-82.88 $\pm$ 5.18
-65.03 $\pm$ 8.51
-44.86 $\pm$ 9.85 8.04 $\pm$ 3.34 24.62 $\pm$ 8.17
-ATM-E (Ours) 83.30 $\pm$ 3.92 65.80 $\pm$ 6.98
-44.84 $\pm$ 7.76
-7.47 $\pm$ 2.84
-23.75 $\pm$ 6.86
 Table 8: Overall accuracy of zero-shot Retrieval on THINGS-EEG dataset.
@@
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet
-76.69 $\pm$ 12.97
-56.98 $\pm$ 16.31
-36.35 $\pm$ 15.11
-5.75 $\pm$ 3.62
-18.14 $\pm$ 9.40
-EEGConformer
-76.17 $\pm$ 13.13
-56.29 $\pm$ 16.70
-34.72 $\pm$ 14.79
-3.98 $\pm$ 2.80
-17.10 $\pm$ 9.21
-ShallowFBCSPNet
-74.32 $\pm$ 12.14
-53.97 $\pm$ 15.81
-33.48 $\pm$ 14.35
-6.10 $\pm$ 4.61
-16.53 $\pm$ 9.94
-EEGNetV4
-92.81 $\pm$ 2.22
-83.15 $\pm$ 4.20
-67.81 $\pm$ 6.11
-19.51 $\pm$ 5.19
-48.99 $\pm$ 6.75
-B.D.
-78.42 $\pm$ 8.81
-58.24 $\pm$ 12.13
-37.97 $\pm$ 11.38
-5.88 $\pm$ 3.49
-18.61 $\pm$ 7.81
-NICE
-92.73 $\pm$ 2.75
-83.26 $\pm$ 5.47
-67.96 $\pm$ 8.31
-19.32 $\pm$ 5.33
-49.26 $\pm$ 9.69
-MLP
-83.09 $\pm$ 2.54
-66.70 $\pm$ 4.16
-45.43 $\pm$ 4.58
-7.23 $\pm$ 1.66
-25.14 $\pm$ 3.66
-ATM-S (Ours) 94.60 $\pm$ 1.93 86.88 $\pm$ 4.20 73.89 $\pm$ 5.93 26.09 $\pm$ 6.96 58.07 $\pm$ 8.16
-ATM-E (Ours)
-92.99 $\pm$ 2.20
-83.81 $\pm$ 4.46
-68.87 $\pm$ 7.27
-22.40 $\pm$ 6.62
-50.59 $\pm$ 9.59
+EEGITNet 76.69 $\pm$ 12.97 56.98 $\pm$ 16.31 36.35 $\pm$ 15.11 5.75 $\pm$ 3.62 18.14 $\pm$ 9.40
+EEGConformer 76.17 $\pm$ 13.13 56.29 $\pm$ 16.70 34.72 $\pm$ 14.79 3.98 $\pm$ 2.80 17.10 $\pm$ 9.21
+ShallowFBCSPNet 74.32 $\pm$ 12.14 53.97 $\pm$ 15.81 33.48 $\pm$ 14.35 6.10 $\pm$ 4.61 16.53 $\pm$ 9.94
+EEGNetV4 92.81 $\pm$ 2.22 83.15 $\pm$ 4.20 67.81 $\pm$ 6.11 19.51 $\pm$ 5.19 48.99 $\pm$ 6.75
+B.D. 78.42 $\pm$ 8.81 58.24 $\pm$ 12.13 37.97 $\pm$ 11.38 5.88 $\pm$ 3.49 18.61 $\pm$ 7.81
+NICE 92.73 $\pm$ 2.75 83.26 $\pm$ 5.47 67.96 $\pm$ 8.31 19.32 $\pm$ 5.33 49.26 $\pm$ 9.69
+MLP 83.09 $\pm$ 2.54 66.70 $\pm$ 4.16 45.43 $\pm$ 4.58 7.23 $\pm$ 1.66 25.14 $\pm$ 3.66
+ATM-S (Ours) 94.92 $\pm$ 1.45 87.91 $\pm$ 3.14 75.37 $\pm$ 5.77 28.64 $\pm$ 6.39 58.47 $\pm$ 8.97
+ATM-E (Ours) 92.99 $\pm$ 2.20 83.81 $\pm$ 4.46 68.87 $\pm$ 7.27 22.40 $\pm$ 6.62 50.59 $\pm$ 9.59
 Subject independent - leave one subject for test (batch size=1024)
 Methods 2-Way 4-Way 10-Way Top-1 Top-5
-EEGITNet
-77.47 $\pm$ 3.75
-55.75 $\pm$ 6.01
-33.60 $\pm$ 6.16
-3.80 $\pm$ 2.01
-14.02 $\pm$ 3.89
-EEGConformer
-68.02 $\pm$ 7.75
-44.24 $\pm$ 9.25
-23.84 $\pm$ 7.65
-1.66 $\pm$ 1.37
-8.62 $\pm$ 4.20
-ShallowFBCSPNet
-76.40 $\pm$ 4.64
-54.81 $\pm$ 7.09
-32.37 $\pm$ 7.28
-2.49 $\pm$ 1.57
-13.24 $\pm$ 5.98
-EEGNetV4
-82.60 $\pm$ 3.17
-64.28 $\pm$ 5.44
-42.24 $\pm$ 6.10
-6.13 $\pm$ 2.40
-21.23 $\pm$ 5.19
-B.D.
-85.30 $\pm$ 6.24
-69.87 $\pm$ 10.09
-50.82 $\pm$ 11.74
-10.57 $\pm$ 5.26
-29.74 $\pm$ 11.28
-NICE
-83.75 $\pm$ 3.21
-65.82 $\pm$ 5.93
-44.41 $\pm$ 6.12
-7.04 $\pm$ 2.83
-23.55 $\pm$ 4.86
-MLP
-80.49 $\pm$ 2.61
-62.09 $\pm$ 3.88
-40.37 $\pm$ 3.71
-5.67 $\pm$ 1.09
-19.97 $\pm$ 3.02
-ATM-S (Ours)
-87.36 $\pm$ 3.97
-72.80 $\pm$ 7.02
-53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
-ATM-E (Ours) 87.41 $\pm$ 3.07 72.85 $\pm$ 5.51
-53.15 $\pm$ 6.52
-11.12 $\pm$ 3.26
-32.61 $\pm$ 6.78
-Table 7: Accuracy of zero-shot Retrieval on THINGS-EEG for odd-numbered subjects(batch size=16).
-Subject dependent - train and test on one subject (batch size=16)
-Method Subject 1 Subject 3 Subject 5 Subject 7 Subject 9
-top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5
-EEGITNet 8.20 19.45 6.95 24.25 6.30 19.50 8.95 19.00 6.10 21.60
-Conformer 4.75 21.30 9.90 34.15 7.55 28.05 12.25 35.35 13.95 34.10
-ShallowFBCSPNet 7.00 20.90 8.45 28.45 7.75 20.80 6.35 27.60 2.80 18.00
-EEGNetV4 10.85 31.10 17.65 41.20 6.75 21.80 14.15 39.05 10.40 30.50
-B.D. 13.05 36.00 14.45 37.85 10.55 32.00 18.65 46.80 14.60 37.55
-NICE 13.25 35.75 23.15 51.40 12.40 33.45 19.45 50.20 17.55 41.70
-MLP 13.00 38.30 13.90 39.60 8.00 24.85 13.00 39.15 12.65 36.20
-ATM-S (Ours) 19.10 49.05 22.95 56.30 16.10 39.95 25.75 55.40 21.50 48.15
-ATM-E (Ours) 20.75 47.35 25.10 56.45 18.50 44.35 24.70 54.10 23.40 52.40
-Subject independent - leave one subject for test (batch size=16)
-Method Subject 1 Subject 3 Subject 5 Subject 7 Subject 9
-top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5
-EEGITNet 3.40 13.75 1.05 10.80 1.45 7.55 4.30 13.65 0.55 5.95
-Conformer 5.45 20.10 2.00 12.75 2.90 12.95 5.50 16.45 1.05 9.90
-ShallowFBCSPNet 4.25 13.30 1.45 7.65 0.90 7.40 3.75 13.75 0.65 5.95
-EEGNetV4 7.20 22.40 4.60 17.80 3.00 15.75 3.80 18.95 3.40 12.70
-B.D. 6.55 21.45 5.45 17.00 3.10 14.75 6.45 24.00 2.60 12.75
-NICE 4.80 20.05 5.80 18.10 5.85 16.95 5.75 21.20 4.95 17.65
-MLP 3.65 16.20 5.95 20.95 5.90 19.20 4.40 14.25 7.45 23.85
-ATM-S (Ours) 8.30 25.50 11.40 31.30 5.95 21.25 4.80 19.45 4.05 13.30
-ATM-E (Ours) 7.80 22.95 13.75 35.70 4.20 14.30 6.40 25.95 4.50 14.90
-Table 8: Accuracy of zero-shot Retrieval on THINGS-EEG for even-numbered subjects(batch size=16).
-Subject dependent - train and test on one subject (batch size=16)
-Method Subject 2 Subject 4 Subject 6 Subject 8 Subject 10 Overall Avg
-top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5
-EEGITNet 6.55 19.05 7.90 23.10 8.10 23.10 9.80 27.35 9.05 24.85 8.08 23.49
-Conformer 8.10 25.80 13.30 37.10 12.75 36.65 15.75 45.10 14.55 37.35 12.89 36.40
-ShallowFBCSPNet 6.15 23.50 10.85 29.10 9.05 26.95 13.15 34.90 11.00 28.55 10.04 28.60
-EEGNetV4 8.80 28.20 14.30 37.45 15.40 42.45 19.40 42.40 15.15 40.85 14.61 38.27
-B.D. 13.35 38.00 17.45 45.00 14.25 40.35 24.50 54.65 22.10 53.45 18.33 46.29
-NICE 11.45 31.05 20.75 48.95 22.25 50.35 29.75 62.40 24.00 57.40 21.64 50.03
-MLP 10.85 33.10 15.20 37.70 12.15 35.05 18.65 47.55 18.30 45.55 15.03 39.79
-ATM-S (Ours) 15.10 41.00 25.75 55.75 20.05 50.10 34.90 67.50 27.15 58.95 24.59 54.66
-ATM-E (Ours) 15.25 44.90 25.75 55.75 23.00 53.95 39.40 70.25 29.60 63.75 26.26 57.36
-Subject independent - leave one subject for test (batch size=16)
-Method Subject 2 Subject 4 Subject 6 Subject 8 Subject 10 Overall Avg
-top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5 top-1 top-5
-EEGITNet 2.65 11.40 3.25 15.35 8.10 23.10 9.80 27.35 9.05 24.85 6.57 20.41
-Conformer 4.65 18.70 4.25 19.60 12.75 36.65 15.75 45.10 14.55 37.35 10.39 31.48
-ShallowFBCSPNet 2.75 13.70 5.55 15.45 9.05 26.95 13.15 34.90 11.00 28.55 8.30 23.91
-EEGNetV4 7.80 24.30 7.35 23.90 15.40 42.45 19.40 42.40 15.15 40.85 13.02 34.78
-B.D. 5.00 16.40 6.65 22.35 14.25 40.35 24.50 54.65 22.10 53.45 14.50 37.44
-NICE 5.55 21.40 8.20 24.90 22.25 50.35 29.75 62.40 24.00 57.40 17.95 43.29
-MLP 5.50 20.65 5.45 18.35 12.15 35.05 18.65 47.55 18.30 45.55 12.01 33.43
-ATM-S (Ours) 7.25 23.65 14.55 39.80 20.05 50.10 34.90 67.50 27.15 58.95 20.78 47.90
-ATM-E (Ours) 8.65 27.60 14.55 39.80 23.00 53.95 39.40 70.25 29.60 63.75 21.62 47.93
-Table 9: Accuracy of zero-shot Retrieval on THINGS-EEG for odd-numbered subjects(batch size=1024).
-Subject dependent - train and test on one subject (batch size=1024)
-Method Subject 1 Subject 3 Subject 5 Subject 7 Subject 9
-Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
-EEGITNet 0.50 3.45 2.40 7.75 5.35 18.25 8.90 24.90 5.50 22.45
-Conformer 0.50 3.20 0.50 6.90 3.40 17.50 7.10 23.80 6.15 22.75
-ShallowFBCSPNet 0.65 3.35 2.30 6.20 3.35 14.55 7.35 23.15 6.80 18.05
-EEGNetV4 11.45 41.45 20.85 48.70 13.75 42.00 19.05 50.30 15.50 40.60
-B.D. 6.70 23.25 6.70 19.75 1.70 8.10 5.85 20.55 6.80 21.05
-NICE 17.65 50.25 22.05 51.20 9.80 27.65 17.15 49.60 19.60 52.55
-MLP 8.40 24.40 5.60 22.65 6.00 25.15 8.35 24.35 5.45 22.25
-ATM-S (Ours) 25.60 60.40 25.00 62.35 12.90 43.00 30.50 61.50 24.35 51.50
-ATM-E (Ours) 18.85 44.10 24.30 56.10 11.95 32.65 24.85 53.75 20.05 47.75
-Subject independent - leave one subject for test (batch size=1024)
-Method Subject 1 Subject 3 Subject 5 Subject 7 Subject 9
-Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
-EEGITNet 4.45 13.95 1.50 12.05 1.00 8.60 4.50 13.35 1.00 8.95
-Conformer 4.45 17.20 0.70 4.75 0.50 5.30 2.40 10.95 0.50 3.35
-ShallowFBCSPNet 2.00 11.25 1.70 9.25 1.00 4.85 3.80 17.30 1.00 5.75
-EEGNetV4 9.45 24.00 8.20 21.90 5.20 18.75 5.45 21.80 5.65 16.40
-B.D. 5.30 16.60 13.45 35.30 8.15 21.30 15.85 37.75 8.10 27.10
-NICE 4.80 20.05 6.75 20.70 4.15 18.55 5.55 24.45 5.10 17.20
-MLP 4.45 12.30 6.75 24.60 4.90 19.55 4.85 17.60 4.45 19.80
-ATM-S (Ours) 10.45 26.75 11.85 33.80 6.95 23.85 16.05 43.50 4.85 22.70
-ATM-E (Ours) 10.45 26.75 13.45 35.30 8.15 24.75 15.85 39.05 7.55 30.65
-Table 10: Accuracy of zero-shot Retrieval on THINGS-EEG for even-numbered subjects(batch size=1024).
-Subject dependent - train and test on one subject (batch size=1024)
-Method Subject 2 Subject 4 Subject 6 Subject 8 Subject 10 Overall Avg
-Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
-EEGITNet 0.75 5.55 6.00 20.35 7.60 20.80 10.55 28.00 9.90 29.95 7.16 20.93
-Conformer 0.50 3.70 3.30 18.10 4.35 21.95 7.50 25.00 6.50 28.15 4.43 19.38
-ShallowFBCSPNet 0.75 4.75 5.25 16.35 7.80 16.95 13.45 32.20 13.25 29.70 8.10 20.19
-EEGNetV4 17.30 43.60 27.20 58.75 20.55 53.25 22.95 56.45 26.50 54.85 22.90 53.38
-B.D. 1.55 6.90 3.65 15.30 3.55 13.05 9.65 27.00 12.70 31.15 6.22 18.68
-NICE 14.35 40.00 20.00 49.90 18.35 49.75 28.05 62.65 26.25 59.10 21.40 52.28
-MLP 6.40 19.45 7.35 27.50 6.75 25.40 7.00 27.20 10.95 33.05 7.69 26.52
-ATM-S (Ours) 22.00 54.50 31.35 60.90 21.30 51.05 38.80 72.00 29.05 63.50 28.50 60.39
-ATM-E (Ours) 17.00 44.85 23.00 51.25 19.40 47.55 35.65 66.75 29.00 61.15 24.81 54.31
-Subject independent - leave one subject for test (batch size=1024)
-Method Subject 2 Subject 4 Subject 6 Subject 8 Subject 10 Overall Avg
-EEGITNet 4.70 15.20 4.80 14.40 5.50 17.45 3.70 14.25 6.90 22.00 5.12 16.66
-Conformer 1.40 9.70 1.15 8.20 1.15 9.70 0.80 5.15 3.55 11.95 1.61 8.94
-ShallowFBCSPNet 2.50 12.15 2.60 17.20 2.15 16.95 1.80 13.15 6.30 24.55 3.07 16.80
-EEGNetV4 8.25 28.15 9.45 26.05 6.40 24.60 9.20 23.60 10.15 36.95 8.69 27.87
-B.D. 4.45 23.20 4.30 12.90 11.15 33.70 18.05 46.30 16.90 43.30 10.97 31.88
-NICE 4.90 22.80 9.65 29.70 8.35 24.35 7.90 25.00 13.25 32.75 8.81 26.92
-MLP 6.50 21.55 9.50 25.15 3.60 16.15 7.25 29.25 12.80 34.35 7.93 25.29
-ATM-S (Ours) 7.10 24.75 14.65 39.40 11.10 35.80 14.95 40.25 20.45 46.50 13.65 37.34
-ATM-E (Ours) 10.45 27.75 14.65 35.25 10.35 35.20 13.95 38.60 15.85 42.85 13.05 35.93
+EEGNetV4 82.85 $\pm$ 3.62 64.65 $\pm$ 6.29 42.35 $\pm$ 7.10 6.25 $\pm$ 2.56 20.95 $\pm$ 5.73
+EEGConformer 56.54 $\pm$ 4.00 31.80 $\pm$ 3.20 13.89 $\pm$ 2.01 0.87 $\pm$ 0.33 4.42 $\pm$ 1.20
+ShallowFBCSPNet 75.76 $\pm$ 2.49 53.63 $\pm$ 3.58 31.43 $\pm$ 4.30 2.51 $\pm$ 1.31 12.03 $\pm$ 2.78
+EEGNetV4 82.60 $\pm$ 3.17 64.28 $\pm$ 5.44 42.24 $\pm$ 6.10 6.13 $\pm$ 2.40 21.23 $\pm$ 5.19
+B.D. 72.84 $\pm$ 12.41 51.41 $\pm$ 15.10 31.67 $\pm$ 12.96 4.10 $\pm$ 2.72 14.46 $\pm$ 7.97
+NICE 83.88 $\pm$ 2.57 66.14 $\pm$ 5.30 45.13 $\pm$ 5.85 8.24 $\pm$ 3.01 23.76 $\pm$ 5.09
+MLP 75.80 $\pm$ 2.45 55.08 $\pm$ 3.07 34.05 $\pm$ 2.83 4.46 $\pm$ 0.81 15.26 $\pm$ 2.34
+ATM-S (Ours) 87.36 $\pm$ 3.97 72.80 $\pm$ 7.02 53.80 $\pm$ 8.41 11.84 $\pm$ 4.80 33.73 $\pm$ 8.73
+ATM-E (Ours) 80.65 $\pm$ 3.11 61.65 $\pm$ 5.31 39.66 $\pm$ 6.44 7.00 $\pm$ 2.20 21.12 $\pm$ 5.25